### v1

# Monte-Carlo Integration Based Multiple-Scattering Channel Modeling for Ultraviolet Communications in Turbulent Atmosphere

Abstract
Modeling of multiple-scattering channels in atmospheric turbulence is essential for the performance analysis of long-distance non-line-of-sight (NLOS) ultraviolet (UV) communications. Existing works on the turbulent channel modeling for NLOS UV communications either ignored the turbulence-induced scattering effect or erroneously estimated the turbulent fluctuation effect, resulting in a contradiction with reported experiments. In this paper, we establish a comprehensive multiple-scattering turbulent channel model for NLOS UV communications considering both the turbulence-induced scattering effect and the turbulent fluctuation effect. We first derive the turbulent scattering coefficient and turbulent phase scattering function based on the Booker-Gordon turbulent power spectral density model. Then an improved estimation method is proposed for both the turbulent fluctuation and the turbulent fading coefficient based on the Monte-Carlo integration approach. Numerical results demonstrate that the turbulence-induced scattering effect can always be ignored for typical UV communication scenarios. Besides, the turbulent fluctuation will increase as either the communication distance, the elevation angle, or the divergence angle increases, which is compatible with existing experimental results. Moreover, we find that the probability density of the equivalent turbulent fading for multiple-scattering turbulent channels can be approximated as a Gaussian distribution.
Index Terms:
Channel modeling, multiple-scattering, turbulent channels, ultraviolet communications
I Introduction
I-A Background and Motivation
The ultraviolet (UV) communication [1, 2] employing the “solar-blind” UV signals ($200$-$280$ nm) as the information carriers has attracted increasing attention in recent decades. Due to its inherent advantages including low background noise, high local security, good adaptivity to extreme weather, and ability of non-line-of-sight (NLOS) links, the UV communication becomes a promising technique for communication scenarios where radio-frequency (RF) communications are impermissible or line-of-sight (LOS) links for traditional free-space optical (FSO) are unavailable [1, 2]. Considerable works have been performed on the NLOS UV channel modeling [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], channel estimation [21, 22, 23, 24, 25, 26], coding and modulation [27, 28, 29, 30], diversity reception [31, 32, 33], duplex and relay techniques [34, 35, 36, 37], and experimental tests [38, 39, 40, 41, 42, 43]. However, most existing works [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] on the channel modeling ignored the impacts of atmospheric turbulence, which cannot be applied to the turbulent cases under either long communication distances or large elevation angles.
The major effects of the atmospheric turbulence on the photon propagating process can be divided into two types: the turbulence-induced scattering effect when a photon interacts with the atmospheric medium [44] and the turbulent fluctuation effect when a photon travels in the atmosphere [45]. In existing works [46, 47, 48, 49, 50, 51, 52], the turbulence-induced scattering effects [46] and the turbulent fluctuation effects [47, 48, 49, 50, 51, 52] were separately studied in the channel modeling of NLOS UV communications. Besides, there exist some contradictions between the existing theoretical turbulent fluctuation model [52] and experimental results [53, 54]. Therefore, a comprehensive turbulent channel model considering both the turbulence-induced scattering effect and the turbulent fluctuation effect is urgently demanded for long-distance NLOS UV communications.
I-B Related Works
The turbulence-induced scattering effect on light wave propagation in the atmosphere was elaborately studied and summarized in [44], where the turbulent atmosphere is modeled as a random continuum with refractive-index varying randomly and continuously in time and space. Some preliminary results on the receiving power and channel impulse response (CIR) of NLOS optical links were derived in [44] based on a single-scattering assumption. The impact of the turbulence-induced scattering effect on the NLOS UV communication was first studied in [46], where the authors claimed that more photons are scattered to the receiver in stronger turbulence for single-scattering cases. However, the impact of the turbulence-induced scattering on the phase scattering function for NLOS UV communications was not clearly studied. Besides, a recent study [55] in underwater optical communication also incorporated the turbulence-induced scattering effect into the scattering-based channel model and demonstrated an addition of more than $5$ dB path loss when the turbulence is considered, which seems opposite to the numerical results obtained in [46]. Therefore, the impact of the turbulence-induced scattering effect on the NLOS UV communication is still not clearly indicated.
The turbulent fluctuation effect on light wave propagating for LOS optical links was elaborately studied and summarized in [45]. The impact of turbulent fluctuation effect on the NLOS UV communications was first studied in a single-scattering model under small common volume assumption [47], where the NLOS link was divided into two LOS links: one from the transmitter to the common volume and another from the common volume to the receiver. When two turbulent fluctuations with log-normal (LN) distribution are introduced into these two LOS links, the authors in [47] demonstrated that the overall turbulent fluctuation can also be approximated as an LN distribution. Based on [47], the bit-error rate performance of NLOS UV communications using this turbulent fluctuation model was studied in [48]. Besides, the turbulent single-scattering model in [47] was extended to narrow beam cases in [49]. A turbulence-induced attenuation effect for LOS links was introduced into the turbulent channel model of NLOS UV communications in [49, 50, 51]. However, the turbulence-induced attenuation considered in LOS laser communications comes from the beam wandering and beam spreading effects [56] caused by the misalignment between the transceivers. Therefore, turbulence-induced attenuation does not apply to NLOS UV links since no alignment is required in NLOS UV communications. Besides, these works [47, 48, 49, 50, 51] are based on the single-scattering model, which cannot be applied to multiple-scattering based long-distance NLOS UV communications.
The turbulent fluctuation model for multiple-scattering was first studied in [52], where the Monte-Carlo based multiple-scattering model [13, 14, 15, 16, 17, 18, 19] is adopted and a Gamma-Gamma (GG) turbulent fading is introduced into each LOS link between two scatters. However, the authors in [52] estimated the turbulent fluctuation by simply calculating the variance of the average receiving power, which cannot really capture the turbulent fluctuation effect. This is because the randomness of the obtained average receiving power in a Monte-Carlo based channel model comes from not only the turbulent fluctuation but also the randomness of the Monte-Carlo process itself [15, 17]. As we will demonstrate in Section IV-C, any attempt to calculate the turbulent fluctuation by estimating the average receiving power of Monte-Carlo based channel models will fail. Based on the same reason, the distribution of the receiving power obtained in [52] cannot characterize the turbulent fading either. Besides, simulation results on the turbulent fluctuation obtained in [52] indicated that the multiple-scattering effect can somehow mitigate the turbulent fluctuation effect, which is opposite to reported experimental results [53, 54]. Therefore, a more sophisticated turbulent fluctuation model is demanded for characterizing the multiple-scattering turbulent channel.
I-C Contributions
Our work aims to establish a multiple-scattering turbulent channel model considering both the turbulence-induced scattering effect and the turbulent fluctuation effect. To achieve this, we first derived the turbulent scattering coefficient and turbulent phase scattering function based on the Booker-Gordon turbulent power spectral density model. Then we proposed an improved estimation method based on a Monte-Carlo integration (MCI) approach for both the turbulent fluctuation and the distribution of the turbulent fading coefficient. The numerical results verified our theoretical analysis and some interesting results were obtained. Specifically, we summarize the main contributions of this work as follows:
  • We established the first comprehensive multiple-scattering turbulent channel model considering both the turbulence-induced scattering effect and fluctuation effect for NLOS UV communications by using an MCI approach.
  • We proposed an improved estimation method for turbulent fluctuation and further proposed an estimation method for the distribution of the turbulent fading coefficient for NLOS UV communications under weak turbulence.
  • We demonstrated that the turbulence-induced scattering effect cannot be ignored only if the turbulent correlation distance approximates the light wavelength under strong turbulence. Besides, for typical UV communication scenarios, we demonstrated that the turbulence-induced scattering effect can always be ignored.
  • We demonstrated that the turbulent fluctuation for NLOS UV communications will increase as either the communication distance, the elevation angle, or the divergence angle increases, which is compatible with existing experimental results.
  • We demonstrated that the distribution of the turbulent fading coefficient for an $n$-order scattering under weak turbulence can be well fitted by a weighted summation of $n$ Gaussian distributions. Besides, the distribution of the overall turbulent fading coefficient for NLOS UV communications can be approximated as a Gaussian distribution.
The rest of this paper is organized as follows. We first introduce the photon propagating model of turbulent channels in Section II. Based on the photon propagating model, we then derive the receiving power of multiple-scattering turbulent channels in Section III. Then we introduce the estimation methods for both the average receiving power, the CIR, the turbulent fluctuation, and the distribution of the turbulent fading coefficient using an MCI approach in Section IV. We then present some numerical results in Section V and conclude our work in Section VI.
II Photon Propagating in Turbulent Channels
The non-turbulent atmosphere can be regarded as a group of particles, which can be modeled as randomly distributed scatters including molecules and aerosols [44]; whereas the turbulent atmosphere is usually regarded as a group of turbulent eddies according to Kolmogorov’s theory of turbulence, which can be modeled as random continuum with refractive-index varying randomly and continuously in time and space [44]. Similar to [55], we consider the scattering effects due to both the random scatterers and the random continuum in this paper. Therefore, the potential scatters include both the particles and the turbulent eddies.
II-A Scattering Effects in Turbulent Channels
II-A1 Scattering Effect of Particles
When the particle size is much smaller than the light wavelength, the scattering process of a light wave can be modeled as a Rayleigh scattering; whereas when the particle size is comparable to or larger than the light wavelength, the scattering process can be modeled as a Mie scattering [1, 2]. We denote the Rayleigh scattering coefficient and the Mie scattering coefficient by $k_{s}^{ray}$ and $k_{s}^{mie}$, respectively. Then the total scattering coefficient due to the randomly distributed particles is given by $k^{par}_{s}=k_{s}^{ray}+k_{s}^{mie}$.
The scattering strength at different directions around the particle is characterized by the phase scattering function [44]. For UV signals, the phase scattering functions for Rayleigh scattering and Mie scattering are respectively given by [1, 2] $p^{ray}(\theta_{s})=\frac{3[1+3\gamma+(1-\gamma)\cos^{2}\theta_{s}]}{16\pi(1+2\gamma)}$ and $p^{mie}(\theta_{s})=\frac{1-g^{2}}{4\pi}\left[\frac{1}{(1+g^{2}-2g\cos\theta_{s})^{{3}/{2}}}\right.$ $\left.+f\frac{(3\cos^{2}\theta_{s}-1)}{2(1+g^{2})^{{3}/{2}}}\right]$, where $\theta_{s}$ is the scattering (zenith) angle between the incident light direction and the scattering light direction; $\gamma$, $g$, and $f$ are the model parameters. The Rayleigh scattering approximates an isotropic scattering and the Mie scattering is a forward direction dominated scattering [1, 2].
II-A2 Scattering Effect of Turbulence
The turbulence-induced scattering effect for a light propagating in a random continuum can be characterized by the differential cross section per unit volume $\sigma(\bm{i},\bm{o})$ [44], where $\bm{i}$ and $\bm{o}$ are the unit vectors for the incident and the scattering directions, respectively. For a statistically homogeneous and isotropic random continuum and a scalar light wave, the differential cross section $\sigma(\bm{i},\bm{o})$ can be obtained as [44]
$\sigma(\bm{i},\bm{o})=2\pi k^{4}\Phi_{n}(k_{s}),$ (1)
where $\Phi_{n}(k_{s})$ is the power spectral density of the turbulence; $k={2\pi}/{\lambda}$ is the wave number of the incident light; $k_{s}=2k\sin({\theta_{s}}/{2})$ and $\theta_{s}$ is the scattering angle between $\bm{i}$ and $\bm{o}$. Then the scattering coefficient caused by the turbulence can be obtained as [44]
$k_{s}^{tur}=\int_{\Omega}\sigma(\bm{i},\bm{o})\mathrm{d}\omega=4\pi^{2}k^{2}\int_{0}^{2k}\Phi_{n}(k_{s})k_{s}\mathrm{d}k_{s},$ (2)
where $\mathrm{d}\omega$ is the differential solid angle and $\Omega$ is the whole solid angle.
From (1) we can observe that the differential cross section $\sigma(\bm{i},\bm{o})$ is a function of the relative direction between $\bm{i}$ and $\bm{o}$. Then we can rewrite $\sigma(\bm{i},\bm{o})$ as $\sigma(\theta_{s})=2\pi k^{4}\Phi_{n}\left(2k\sin({\theta_{s}}/{2})\right)$ and further define the phase scattering function due to the turbulence as
$p^{tur}(\theta_{s})$ $\triangleq\frac{\sigma(\theta_{s})}{\int_{\Omega}\sigma(\theta_{s})\mathrm{d}\omega}=\frac{2\pi k^{4}}{k_{s}^{tur}}\Phi_{n}\left(2k\sin({\theta_{s}}/{2})\right).$ (3)
The scattering coefficient $k_{s}^{tur}$ in (2) and the phase scattering function $p^{tur}(\theta_{s})$ in (3) are closely related to the power spectral density $\Phi_{n}(k_{s})$ of the turbulence. However, different assumptions on the turbulence can result in different forms of the power spectral density [44]. Without loss of generality, we adopt the widely used Booker-Gordon model in this paper [44, 46].
The Booker-Gordon model characterizes the turbulent medium by using two quantities, i.e., the variance of the refractive index $\langle n_{1}^{2}\rangle$ and the correlation distance $d_{0}$. Here $n_{1}\triangleq n-\langle n\rangle$ is the fluctuation of the refractive index $n$; $d_{0}$ represents the average size of the turbulent eddies. Then the correlation function of the refractive-index at two positions $\bm{r}_{1}$ and $\bm{r}_{2}$ in Booker-Gordon model is assumed to be [44, 46] $B_{n}(r_{d})\triangleq\langle n_{1}(\bm{r}_{1})n_{1}(\bm{r}_{2})\rangle=\langle n_{1}^{2}\rangle\exp\left({-{r_{d}}/{d_{0}}}\right)$, where $r_{d}=|\bm{r}_{1}-\bm{r}_{2}|$. According to the Wiener-Khinchin theorem [44, 46], the power spectral density $\Phi_{n}(k_{s})$ is the Fourier transform of the correlation function $B_{n}(r_{d})$. For a scalar wave, the power spectral density $\Phi_{n}(k_{s})$ can be obtained as [44, 46]
$\Phi_{n}(k_{s})$ $=\frac{\langle n_{1}^{2}\rangle d_{0}^{3}}{\pi^{2}(1+k_{s}^{2}d_{0}^{2})^{2}}.$ (4)
Then by substituting (4) into (2), we can derive the scattering coefficient of Booker-Gordon model as
$k_{s}^{tur}$ $=\frac{8k^{4}\langle n_{1}^{2}\rangle d_{0}^{3}}{1+4k^{2}d_{0}^{2}}.$ (5)
Besides, when the turbulence is assumed to be isotropic, we have [44, 46] $\langle n_{1}^{2}\rangle=\frac{C_{n}^{2}L_{0}^{{2}/{3}}}{1.91}$, where $L_{0}$ is the outer scale parameter representing the largest turbulent eddy size; $C_{n}^{2}$ is the refractive-index structure parameter representing the strength of the turbulence. The refractive-index structure parameter $C_{n}^{2}$ is defined as the constant coefficient of the refractive-index structure equation [44] $D_{n}(r)\triangleq\langle[n(\bm{r}_{1})-n(\bm{r}_{2})]^{2}\rangle=C_{n}^{2}r^{{2}/{3}}$, where $n(\bm{r}_{i})$ is the refractive-index at position $\bm{r}_{i}$ and $r=|\bm{r}_{1}-\bm{r}_{2}|$ is the distance between two positions $\bm{r}_{1}$ and $\bm{r}_{2}$. The refractive-index structure equation also provides a way for measuring $C_{n}^{2}$ in practical implementations. Typical values of $L_{0}$ can vary from the magnitude of $10^{0}$ m to $10^{3}$ m; typical values of $C_{n}^{2}$ are at the magnitude of $10^{-15}$ $\text{m}^{-2/3}$ and $C_{n}^{2}$ can vary from $10^{-17}$ for weak turbulence to $10^{-13}$ for strong turbulence [44, 46]. Then by substituting $\langle n_{1}^{2}\rangle$ into (5), we can further obtain
$k_{s}^{tur}=\frac{8k^{4}C_{n}^{2}L_{0}^{{2}/{3}}d_{0}^{3}}{1.91(1+4k^{2}d_{0}^{2})}.$ (6)
From the expression of $k_{s}^{tur}$ in (6), we can observe that $k_{s}^{tur}\propto C_{n}^{2}$. Besides, when $k^{2}d_{0}^{2}\ll 1$, we have $k_{s}^{tur}\approx\frac{8k^{4}C_{n}^{2}L_{0}^{{2}/{3}}d_{0}^{3}}{1.91}$, which indicates that $k_{s}^{tur}\propto d_{0}^{3}$ when the correlation distance $d_{0}$ is much smaller than the wavelength $\lambda$. In this case, we also have $k_{s}^{tur}\propto 1/\lambda^{4}$, which coincides with the Rayleigh scattering. When $k^{2}d_{0}^{2}\gg 1$, we have $k_{s}^{tur}\approx\frac{2k^{2}C_{n}^{2}L_{0}^{{2}/{3}}d_{0}}{1.91}$, which indicates that $k_{s}^{tur}\propto d_{0}$ when the correlation distance $d_{0}$ is much larger than the wavelength $\lambda$.
Substituting (4) and (5) into (3), we can obtain the phase scattering function for the Booker-Gordon model as
$p^{tur}(\theta_{s})=\frac{1+4k^{2}d_{0}^{2}}{4\pi(1+4k^{2}d_{0}^{2}\sin^{2}({\theta_{s}}/{2}))^{2}}.$ (7)
From the phase scattering function $p^{tur}(\theta_{s})$ in (7), we can observe that the phase scattering function depends only on the wave number $k$ and the correlation distance $d_{0}$. Besides, when $k^{2}d_{0}^{2}\ll 1$, we have $p^{tur}(\theta_{s})\approx 1/4\pi$, which indicates that the scattering will approach an isotropic scattering when the correlation distance $d_{0}$ is much smaller than the wavelength $\lambda$. This isotropic scattering property is similar to the Rayleigh scattering. When $k^{2}d_{0}^{2}\gg 1$, the majority of the scattering intensity happens at small scattering angles. This indicates that the scattering is forward direction dominated when the correlation distance $d_{0}$ is much larger than the wavelength $\lambda$, which is similar to the Mie scattering.
Combining the coefficients $k_{s}^{ray}$, $k_{s}^{mie}$, and $k_{s}^{tur}$, we can obtain the total scattering coefficients as
$k_{s}^{tot}=k_{s}^{ray}+k_{s}^{mie}+k_{s}^{tur}.$ (8)
Similarly, combining the phase scattering functions $p^{ray}(\theta_{s})$, $p^{mie}(\theta_{s})$, and $p^{tur}(\theta_{s})$, we can obtain the total phase scattering function as a weighted function of $p^{ray}(\theta_{s})$, $p^{mie}(\theta_{s})$, and $p^{tur}(\theta_{s})$, i.e.,
$p^{tot}(\theta_{s})=\frac{k_{s}^{ray}}{k_{s}^{tot}}p^{ray}(\theta_{s})+\frac{k_{s}^{mie}}{k_{s}^{tot}}p^{mie}(\theta_{s})+\frac{k_{s}^{tur}}{k_{s}^{tot}}p^{tur}(\theta_{s}).$ (9)
II-B Absorption Effects in Turbulent Channels
[…]
Denoting the absorption coefficient due to the particles by $k_{a}^{par}$, we can obtain the total extinction coefficient of the atmosphere as
$k^{tot}_{e}\triangleq k_{s}^{tot}+k^{par}_{a}=k_{s}^{ray}+k_{s}^{mie}+k_{s}^{tur}+k^{par}_{a}.$ (10)
II-C Fluctuation Effects in Turbulent Channels
Besides the scattering and absorption effects, the turbulence also introduces a fluctuation effect on the light irradiance [45]. The turbulent fluctuation effect can be characterized by the turbulent fading coefficient $\eta$ with $\eta\triangleq P_{r}/P_{t}$ and $\quad\eta\geq 0$, where $P_{t}$ and $P_{r}$ are the transmit power and receiving power of a traveling process, respectively.
For weak turbulent conditions, the turbulent fading coefficient is usually modeled as an LN distribution with the normalized probability density function (PDF) given by [45]
$f_{H,LN}(\eta;d)=\frac{1}{\eta\sqrt{2\pi\sigma_{I}^{2}}}\exp\left({-\frac{(\ln\eta+0.5\sigma_{I}^{2}(d))^{2}}{2\sigma_{I}^{2}(d)}}\right),$ (11)
where $\sigma_{I}^{2}(d)=e^{\sigma_{r}^{2}(d)}-1$ and $\sigma_{r}^{2}(d)$ is the Rytov variance for a propagating distance $d$. For a plane wave, the Rytov variance is given by [45]
$\sigma_{r}^{2}(d)=1.23C_{n}^{2}k^{7/6}d^{11/6}.$ (12)
For moderate and strong turbulent conditions, the turbulent fading coefficient is usually modeled as a GG distribution with the normalized PDF given by [45, 57]
$f_{H,GG}(\eta;d)$ $=\frac{2[\alpha(d)\beta(d)]^{\frac{\alpha(d)+\beta(d)}{2}}\eta^{\frac{\alpha(d)+\beta(d)}{2}-1}}{\Gamma(\alpha(d))\Gamma(\beta(d))}K_{\alpha(d)-\beta(d)}\left(2\sqrt{\alpha(d)\beta(d)\eta}\right),\quad\eta\geq 0,$ (13)
where $K_{v}(x)$ is the modified Bessel function of the second kind with parameter $v$; $\alpha(d)$ and $\beta(d)$ are related to the Rytov variance $\sigma_{r}^{2}(d)$ as [45, 57]
$\begin{cases}\alpha(d)=\left[\exp\left(\frac{0.49\sigma_{r}^{2}(d)}{(1+1.11\sigma_{r}^{12/5})^{7/6}}\right)-1\right]^{-1},\\ \beta(d)=\left[\exp\left(\frac{0.51\sigma_{r}^{2}(d)}{(1+0.69\sigma_{r}^{12/5})^{5/6}}\right)-1\right]^{-1}.\end{cases}$ (14)
For weak turbulence, i.e., $\sigma_{r}^{2}\to 0$, we have $\alpha\approx 1/0.49\sigma_{r}^{2}$, $\beta\approx 1/0.51\sigma_{r}^{2}$, and the turbulent variance $\text{Var}[\eta]=1/\alpha+1/\beta+1/\alpha\beta\approx\sigma_{r}^{2}$. For saturated turbulence, i.e., $\sigma_{r}^{2}\to\infty$, we have $\alpha\to\infty$, $\beta\to 1$, and the turbulent variance $\text{Var}[\eta]\approx 1$.
III Receiving Power of Multiple-Scattering in Turbulent Channels
Figure 1: Geometry setting for the multiple-scattering process of NLOS UV communications
We first introduce the geometry setting for the multiple-scattering process of NLOS UV communications, shown in Fig. 1. The receiver (Rx) locates at the origin $(0,0,0)$ and the transmitter (Tx) locates at the $y$-axis with coordinates $(0,r,0)$. The direction cosines of Tx and Rx pointing directions are denoted by $\bm{\mu}_{T}$ and $\bm{\mu}_{R}$, respectively; The zenith angle and the azimuth angle of the Tx(Rx) are denoted by $\theta_{T(R)}$ and $\phi_{T(R)}$, respectively. The divergence angle of the Tx light beam is denoted by $\beta_{T}$ and the divergence angle of the Rx field-of-view (FOV) is denoted by $\beta_{R}$. We regard the initial photon emitting as the $zero$-order scattering and denote the propagating distance, scattering zenith angle and scattering azimuth angle of $i$-order scattering by $d_{i}$, $\theta_{i}$ and $\phi_{i}$, respectively. The position of the $i$th scatter is denoted by $\bm{r}_{i}$. The direction cosine of $i$-order scattering is denoted by $\bm{\mu}_{i}$. The initial scattering zenith angle $\theta_{0}$ is the angle between $\bm{\mu}_{0}$ and $\bm{\mu}_{T}$. We consider an $n$-order scattering process and define $\bm{d}\triangleq[d_{0},d_{1},\cdots,d_{n-1}]^{\text{T}}$, $\bm{\theta}\triangleq[\theta_{0},\theta_{1},\cdots,\theta_{n-1}]^{\text{T}}$, and $\bm{\phi}\triangleq[\phi_{0},\phi_{1},\cdots,\phi_{n-1}]^{\text{T}}$.
III-A Receiving Power Ignoring Turbulent Fluctuation
A photon propagating path of an $n$-order scattering is full characterized by the parameters $\{\bm{d},\bm{\theta},\bm{\phi}\}$. After scattered by the last scatter, the photon can be detected if and only if the position of the $n$th scatter locates in the FOV of the receiver. We denote the set of parameters satisfying this geometrical constrain by $\mathcal{F}\triangleq\{\bm{d},\bm{\theta},\bm{\phi}|\bm{r}_{n}\text{ in FOV}\}=\{\bm{d},\bm{\theta},\bm{\phi}|\bm{r}_{n}\cdot\bm{\mu}_{R}\geq d_{n}\cos(\beta_{R}/2)\}$. Then the conditional receiving probability of an $n$-order scattering process for a given propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$ can be approximated as [17]
$P_{n}(\bm{d},\bm{\theta},\bm{\phi})$ $=I_{n}(\bm{d},\bm{\theta},\bm{\phi})\left(\frac{k_{s}^{tot}}{k_{e}^{tot}}\right)^{n}\exp({-k_{e}^{tot}d_{n}})\cos\phi_{r}\min\left(1,p^{tot}(\theta_{r,n})\Omega_{r}\right),$ (15)
where $d_{n}\triangleq\|\bm{r}_{n}\|$ and $\theta_{r,n}$ is the angle between $\bm{\mu}_{n-1}$ and $-\bm{r}_{n}$; $\phi_{r}$ is the angle between $\bm{\mu}_{R}$ and $\bm{r}_{n}$; $\Omega_{r}$ is the solid angle formed by $\bm{r}_{n}$ and the receiving area $A_{r}$; and we have $\Omega_{r}\approx 2\pi(1-d_{n}/\sqrt{d_{n}^{2}+r_{A}^{2}})$, where $r_{A}=\sqrt{A_{r}/\pi}$ is the radius of the receiving area; $I_{n}(\bm{d},\bm{\theta},\bm{\phi})$ is an indicator function; and we have $I_{n}(\bm{d},\bm{\theta},\bm{\phi})=1$ and $I_{n}(\bm{d},\bm{\theta},\bm{\phi})=0$ when the $n$th scatter locates within and without the FOV, respectively. Then the conditional receiving probability can be rewritten as
$P_{n}(\bm{d},\bm{\theta},\bm{\phi})=\begin{cases}P_{d,n}(\bm{d},\bm{\theta},\bm{\phi}),\quad\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F},\\ 0,\quad\{\bm{d},\bm{\theta},\bm{\phi}\}\notin\mathcal{F},\end{cases}$ (16)
where $P_{d,n}(\bm{d},\bm{\theta},\bm{\phi})\triangleq\left(\frac{k_{s}^{tot}}{k_{e}^{tot}}\right)^{n}\exp({-k_{e}^{tot}d_{n}})\cos\phi_{r}\min\left(1,\right.$ $\left.p^{tot}(\theta_{r,n})\Omega_{r}\right)$.
Then the receiving probability can be obtained by averaging $\{\bm{d},\bm{\theta},\bm{\phi}\}$ out, i.e.,
$P_{n}$ $=\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}P_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi},$ (17)
where $f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})=\prod_{i=0}^{n-1}f_{D}(d_{i})f_{\Theta}(\theta_{i})f_{\Phi}(\phi_{i})$ is the joint PDF of $\{\bm{d},\bm{\theta},\bm{\phi}\}$. We have used the fact that $\bm{d}$, $\bm{\theta}$, and $\bm{\phi}$ are independent from each other. For the propagating distance $d_{i}$ with $i\in\{0,1,\cdots,n-1\}$, we have $f_{D}(d_{i})=k_{e}^{tot}e^{-k_{e}^{tot}d_{i}}$ with $0\leq d_{i}\leq\infty$ [17]. For the scattering angle $\theta_{i}$ with $i=0$, we have $f_{\Theta}(\theta_{0})=\frac{\sin\theta_{0}}{1-\cos({\beta_{T}}/{2})}$ with $0\leq\theta_{0}\leq{\beta_{T}}/{2}$ when a uniform light source is assumed [17]; and for the scattering angle $\theta_{i}$ with $i\in\{1,2,\cdots,n-1\}$, we have $f_{\Theta}(\theta_{i})=2\pi p^{tot}(\theta_{i})\sin\theta_{i}$ with $0\leq\theta_{i}\leq\pi$ [17]. For the scattering azimuth angle $\phi_{i}$ with $i\in\{0,1,\cdots,n-1\}$, we have $f_{\Phi}(\phi_{i})=\frac{1}{2\pi}$ with $0\leq\phi_{i}\leq 2\pi$ [17].
Without loss of generality, we assume that a unit light power is transmitted from the light source. Then the average receiving power equals the receiving probability $P_{n}$.
III-B Receiving Power Considering Turbulent Fluctuation
When the turbulent fluctuation effect is considered, a turbulent fading coefficient is introduced to each photon propagating distance. We denote the turbulent fading coefficient for the $i$th propagating distance $d_{i}$ by $\eta_{i}$. Then a photon propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$ forms $(n+1)$ propagating distances $\{d_{0},d_{1},\cdots,d_{n}\}$, which corresponds to $(n+1)$ turbulent fading coefficients $\{\eta_{0},\eta_{1},\cdots,\eta_{n}\}$.
The conditional receiving probability without turbulent fluctuation for a given propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$ is $P_{n}(\bm{d},\bm{\theta},\bm{\phi})$ given in (15). Then the conditional receiving probability under turbulent fluctuation effect for a given set of turbulent fading coefficients $\bm{\eta}\triangleq[\eta_{0},\eta_{1},\cdots,\eta_{n}]^{\text{T}}$ and a propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$ can be expressed as
$P_{tur,n}(\bm{\eta},\bm{d},\bm{\theta},\bm{\phi})=\prod_{i=0}^{n}\eta_{i}P_{n}(\bm{d},\bm{\theta},\bm{\phi}).$ (18)
Then the receiving probability under turbulent fluctuation for a given propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$ can be obtained by averaging $\bm{\eta}$ out, i.e.,
$P_{tur,n}(\bm{d},\bm{\theta},\bm{\phi})$ $=\int_{\bm{\eta}}P_{tur,n}(\bm{\eta},\bm{d},\bm{\theta},\bm{\phi})f_{H|D,\Theta,\Phi}(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{\eta},$ (19)
where $f_{H|D,\Theta,\Phi}(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})$ is the joint PDF of $\{\eta_{0},\eta_{1},\cdots,\eta_{n}\}$ conditioned on a given propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$. For a given propagating path, $\eta_{0},\eta_{1},\cdots,\eta_{n}$ are assumed as independent variables. Then we have
$f_{H|D,\Theta,\Phi}(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})=\prod_{i=0}^{n}f_{H}(\eta_{i};d_{i}),$ (20)
where $f_{H}(\eta;d)=f_{H,LN}(\eta;d)$ is given in (11) for LN distribution and $f_{H}(\eta;d)=f_{H,GG}(\eta;d)$ is given in (13) for GG distribution.
By substituting (18) and (20) into (19), we can obtain
$P_{tur,n}(\bm{d},\bm{\theta},\bm{\phi})$ $=P_{n}(\bm{d},\bm{\theta},\bm{\phi})\prod_{i=0}^{n}\left[\int_{\eta_{i}}\eta_{i}f_{H}(\eta_{i};d_{i})\mathrm{d}\eta_{i}\right]=P_{n}(\bm{d},\bm{\theta},\bm{\phi}),$ (21)
where we have used the property that $\int_{\eta_{i}}\eta_{i}f_{H}(\eta_{i};d_{i})\mathrm{d}\eta_{i}=1$ for $i=0,1,\cdots,n$.
Then the receiving probability considering the turbulent fluctuation effect can be obtained by averaging $\{\bm{d},\bm{\theta},\bm{\phi}\}$ out, i.e.,
$P_{tur,n}$ $=\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}P_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}=P_{n}.$ (22)
From (22), we can see that $P_{tur,n}=P_{n}$, which indicates that the average receiving power considering the turbulent fluctuation equals the average receiving power ignoring the turbulent fluctuation. Therefore, in practical implementations, we can estimate the average receiving power by solving the integration in (17).
IV Estimating Turbulent Channels Using MCI Approach
IV-A Estimating the Average Receiving Power
The integration in (17) can be estimated by using the Monte-Carlo methods. In this paper, we use the MCI approach to estimate the average receiving power in (17). Compared with another widely used Monte-Carlo simulation (MCS) approach [13, 14], the MCI approach [15, 16, 17, 18, 19] enjoys a more flexible sampling function, resulting in a higher computational efficiency. Specifically, we can rewrite (17) as
$P_{n}=\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}g_{n}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi},$ (23)
where $g_{n}(\bm{d},\bm{\theta},\bm{\phi})$ is defined as $g_{n}(\bm{d},\bm{\theta},\bm{\phi})\triangleq P_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})$, and where $P_{n}(\bm{d},\bm{\theta},\bm{\phi})$ is given in (16). Now we can choose a sampling PDF $f_{n}(\bm{d},\bm{\theta},\bm{\phi})$ and further rewrite the receiving power $P_{n}$ in (23) as
$P_{n}$ $=\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}O_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{n}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}=\text{E}[O_{n}(\bm{d},\bm{\theta},\bm{\phi})],$ (24)
where $O_{n}(\bm{d},\bm{\theta},\bm{\phi})\triangleq g_{n}(\bm{d},\bm{\theta},\bm{\phi})/f_{tur,n}(\bm{d},\bm{\theta},\bm{\phi})$ is called the objective function of MCI approach; and $\text{E}[x]$ denotes the expectation of $x$. Therefore, we have expressed the receiving power $P_{n}$ as the expectation of the objective function $O_{n}(\bm{d},\bm{\theta},\bm{\phi})$ when the joint PDF of $\{\bm{d},\bm{\theta},\bm{\phi}\}$ is the sampling PDF $f_{n}(\bm{d},\bm{\theta},\bm{\phi})$.
Different sampling PDF $f_{n}(\bm{d},\bm{\theta},\bm{\phi})$ can result in different computational efficiency of MCI approach [15, 16, 17]. Here we adopt the so-called important sampling method [15], which enjoys the same convergent speed as the MCS approach. The sampling PDF of the important sampling is given by
$f_{n}(\bm{d},\bm{\theta},\bm{\phi})=f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})=\prod_{i=0}^{n-1}f_{D}(d_{i})f_{\Theta}(\theta_{i})f_{\Phi}(\phi_{i}).$ (25)
According to the sampling PDF $f_{n}(\bm{d},\bm{\theta},\bm{\phi})$ in (25), we can obtain the sampling functions for generating the propagating distance, the scattering angle, and the scattering azimuth angle as
$d_{i}=-\frac{\ln(1-\text{rand}(1))}{k_{e}^{tot}},\quad i=0,1,\cdots,n-1,$ (26)
$\theta_{i}=\begin{cases}\arccos\left(1-\text{rand}(1)\times\left(1-\cos(\beta_{T}/2)\right)\right),\quad i=0,\\ F_{\Theta}^{-1}(\text{rand}(1)),\quad i=1,2,\cdots,n-1,\end{cases}$ (27)
$\phi_{i}=2\pi\times\text{rand}(1),\quad i=0,1,\cdots,n-1,$ (28)
where $\text{rand}(1)$ is a random number between 0 and 1; $F_{\Theta}^{-1}(\theta_{s})$ is the inverse function of the cumulative distribution function of the scattering angle, i.e., $F_{\Theta}(\theta_{s})=\int_{0}^{\theta_{s}}f_{\Theta}(\theta)\mathrm{d}\theta$, which can be obtained as
$F_{\Theta}(\theta_{s})$ $=C_{1}\cos\theta_{s}+C_{2}\cos^{3}\theta_{s}+C_{3}+C_{4}(C_{5}-C_{6}\cos\theta_{s})^{-1/2}+\frac{C_{7}(1-\cos\theta_{s})}{2+C_{8}(1-\cos\theta_{s})},$ (29)
where $C_{1}=-\frac{3(1+3\gamma)}{8(1+2\gamma)}\frac{k_{s}^{ray}}{k_{s}^{tot}}+\frac{f(1-g^{2})}{4(1+g^{2})^{3/2}}\frac{k_{s}^{mie}}{k_{s}^{tot}}$, $C_{2}=-\frac{1-\gamma}{8(1+2\gamma)}\frac{k_{s}^{ray}}{k_{s}^{tot}}-\frac{f(1-g^{2})}{4(1+g^{2})^{3/2}}\frac{k_{s}^{mie}}{k_{s}^{tot}}$, $C_{3}=\frac{1}{2}\frac{k_{s}^{ray}}{k_{s}^{tot}}+\frac{1+g}{2g}\frac{k_{s}^{mie}}{k_{s}^{tot}}$, $C_{4}=-\frac{1-g^{2}}{2g}\frac{k_{s}^{mie}}{k_{s}^{tot}}$, $C_{5}=1+g^{2}$, $C_{6}=2g$, $C_{7}=(1+4k^{2}d_{0}^{2})\frac{k_{s}^{tur}}{k_{s}^{tot}}$, and $C_{8}=4k^{2}d_{0}^{2}$. For $i=1,2,\cdots,n-1$, the scattering angle $\theta_{i}$ can be solved by using numerical methods, e.g., Newton’s bisection search.
Then the corresponding objective function becomes
$O_{n}(\bm{d},\bm{\theta},\bm{\phi})$ $=P_{n}(\bm{d},\bm{\theta},\bm{\phi})=\begin{cases}P_{d,n},\quad\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F},\\ 0,\quad\{\bm{d},\bm{\theta},\bm{\phi}\}\notin\mathcal{F}.\end{cases}$ (30)
According to the law of large numbers, we can estimate the expectation $P_{n}$ using the average of sampling values for the objective function. Specifically, we can first randomly generate $M$ sampling points $\{\bm{s}_{n}^{1},\bm{s}_{n}^{2},\cdots,\bm{s}_{n}^{M}\}$ according to the sampling PDF given in (25), where we have
$\bm{s}_{n}^{m}\triangleq[d_{0}^{m},\theta_{0}^{m},\phi_{0}^{m},\cdots,d_{n-1}^{m},\theta_{n-1}^{m},\phi_{n-1}^{m}]^{\text{T}}.$ (31)
Then we can approximate the receiving power when $M$ is large as
$P_{n}=\frac{1}{M}\sum_{m=1}^{M}O_{n}(\bm{s}_{n}^{m}).$ (32)
The total receiving power over $N$ scattering orders can be obtained as
$P_{tot}=\sum_{n=1}^{N}P_{n}=\frac{1}{M}\sum_{m=1}^{M}\sum_{n=1}^{N}O_{n}(\bm{s}_{n}^{m}).$ (33)
IV-B Estimating the Channel Impulse Response
For each sampling point $\bm{s}_{n}^{m}$, we can calculating the corresponding receiving time as
$t_{n}^{m}=\frac{\sum_{i=0}^{n}d_{i}^{m}}{c},$ (34)
where $c$ is the light speed in the atmosphere. Then according to the result in [17], we can obtain the CIR of the multiple-scattering of UV communications in the following steps:
  • Step 1: Classify the objective function value $O_{n}(\bm{s}_{n}^{m})$ into $L$ different time slots $\{(0,\Delta t)$, $(\Delta t,2\Delta t),\cdots,((L-1)\Delta t,L\Delta t)\}$ according to its receiving time $t_{n}^{m}$, where $\Delta t$ is a pre-chosen time interval;
  • Step 2: Add up objective function values in each time slot and divide it by the number of sampling points $M$. Then we obtain the receiving probability sequences $[{P}_{n}(\Delta t),{P}_{n}(2\Delta t)$, $\cdots,{P}_{n}(L\Delta t)]^{\text{T}}$;
  • Step 3: The CIR is obtained by normalizing the sequences $[{P}_{n}(\Delta t),{P}_{n}(2\Delta t),\!\cdots,\!{P}_{n}(L\Delta t)]^{\text{T}}$ using the receiving time interval $\Delta t$ and the receiving area $A_{r}$. Therefore, the value of CIR in a unit of irradiance at $t=l\Delta t$ is given by
    $h_{n}(l\Delta t)=\frac{P_{n}(l\Delta t)}{\Delta_{t}A_{r}},\quad l=1,2\cdots,L.$ (35)
Then the total CIR at $t=l\Delta t$ over $N$ scattering orders can be obtained as
$h_{tot}(l\Delta t)=\sum_{n=1}^{N}\frac{P_{n}(l\Delta t)}{\Delta_{T}A_{r}},\quad l=1,2\cdots,L.$ (36)
IV-C Estimating the Turbulent Fluctuation
It is challenging to estimate the turbulent fluctuation of the receiving power by using Monte-Carlo based methods. This is because the Monte-Carlo process itself will introduce a fluctuation effect on the calculated average receiving power [17]. In existing literature [52], the turbulent fluctuation was estimated by simply calculating the variance of the average receiving power $P_{tur,n}$, i.e, $\text{Var}[P_{tur,n}]$, obtained from the Monte-Carlo models. However, because $P_{tur,n}=P_{n}$, using the expression in (32), we can obtain
$\text{Var}[P_{tur,n}]=\frac{1}{M^{2}}\sum_{m=1}^{M}\text{Var}[O_{n}(\bm{s}_{n}^{m})]=\frac{1}{M}\text{Var}[O_{n}(\bm{d},\bm{\theta},\bm{\phi})],$ (37)
where we have used the fact that all sampling points $\{\bm{s}_{n}^{1},\bm{s}_{n}^{2},\cdots,\bm{s}_{n}^{M}\}$ are independent from each other and $\text{Var}[O_{n}(\bm{s}_{n}^{1})]=\text{Var}[O_{n}(\bm{s}_{n}^{2})]\cdots=\text{Var}[O_{n}(\bm{s}_{n}^{M})]=\text{Var}[O_{n}(\bm{d},\bm{\theta},\bm{\phi})]$ is the variance of the objective function. We can see that the variance $\text{Var}[P_{tur,n}]$ depends on both $\text{Var}[O_{n}(\bm{d},\bm{\theta},\bm{\phi})]$ and $M$. Besides, we have $\text{Var}[P_{tur,n}]\to 0$ when the number of sampling points $M\to\infty$, which means the calculated variance can be arbitrarily small as long as $M$ is large enough. Whereas, larger sampling points $M$ in the Monte-Carlo process only reduce the fluctuation caused by randomness, leading to more stable and precise results. Therefore, any attempt to calculate the turbulent fluctuation by estimating $\text{Var}[P_{tur,n}]$ will fail due to the fact that $\text{Var}[P_{tur,n}]$ depends on $M$. Here we present an improved estimating method for the turbulent fluctuation, which is independent of $M$.
By substituting (16) into (18), we can obtain
$P_{tur,n}(\bm{\eta},\bm{d},\bm{\theta},\bm{\phi})=\begin{cases}\prod_{i=0}^{n}\eta_{i}P_{d,n}(\bm{d},\bm{\theta},\bm{\phi}),\quad\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F},\\ 0,\quad\{\bm{d},\bm{\theta},\bm{\phi}\}\notin\mathcal{F},\end{cases}$ (38)
which can be regarded as a function of random variables $\bm{\eta}$, $\bm{d}$, $\bm{\theta}$, and $\bm{\phi}$. Therefore, the randomness of the receiving power comes from both the turbulent fading $\bm{\eta}$ and the propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$. To characterize the turbulent fluctuation effect only, we can average $\{\bm{d},\bm{\theta},\bm{\phi}\}$ out and obtain the instantaneous receiving power conditioned on turbulent fading coefficients as
$P_{tur,n}(\bm{\eta})$ $=\!\prod_{i=0}^{n}\eta_{i}\!\!\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}\!P_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}=\!{\prod_{i=0}^{n}\eta_{i}}P_{n}.$ (39)
Then we can define an equivalent turbulent fading coefficient $\eta_{eq,n}\triangleq P_{tur,n}(\bm{\eta})/P_{n}$ on the average receiving power $P_{n}$ such that the instantaneous receiving power $P_{tur,n}(\bm{\eta})=\eta_{eq,n}P_{n}$. Obviously, we have
$\eta_{eq,n}=\prod_{i=0}^{n}\eta_{i}.$ (40)
One can easily verify that $\text{E}[\eta_{eq,n}]=1$. Then the turbulent fluctuation for the $n$th scattering order can be fully characterized by the turbulent variance
$\sigma_{tur,n}^{2}$ $\triangleq\text{Var}[\eta_{eq,n}]=\int_{\bm{\eta}}(\eta_{eq,n}-1)^{2}f_{H}(\bm{\eta})\mathrm{d}\bm{\eta},$ (41)
where $f_{H}(\bm{\eta})$ is the joint PDF of $\{\eta_{0},\eta_{1},\cdots,\eta_{n}\}$. From (38) we can see that those propagating paths with $\{\bm{d},\bm{\theta},\bm{\phi}\}\notin\mathcal{F}$ contribute zero to the turbulent fluctuation effect. Therefore, we should restrict the variance on the set $\mathcal{F}$. Then the joint PDF $f_{H}(\bm{\eta})$ can be obtained by averaging $\{\bm{d},\bm{\theta},\bm{\phi}\}$ out of the conditional PDF $f_{H|D,\Theta,\Phi}(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})$ in (20) on the set $\mathcal{F}$, i.e.,
$f_{H}(\bm{\eta})$ $\!=\!\int_{\mathcal{F}}f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi}|\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F})f_{H|D,\Theta,\Phi}(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi},$ (42)
where $f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi}|\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F})$ is the joint PDF of $\{\bm{d},\bm{\theta},\bm{\phi}\}$ on the set $\mathcal{F}$; and we have
$f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi}|\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F})$ $=\frac{f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})}{\text{Pr}[\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F}]},$ (43)
where $\text{Pr}[\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F}]$ is the probability of $\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F}$ given by
$\text{Pr}[\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F}]=\int_{\mathcal{F}}f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}.$ (44)
By substituting (20), (43), and (44) into (42), we can obtain
$f_{H}(\bm{\eta})$ $=\frac{\int_{\mathcal{F}}\prod_{i=0}^{n}f_{H}(\eta_{i};d_{i})f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\mathcal{F}}f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}$ (45)
$=\frac{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}\prod_{i=0}^{n}f_{H}(\eta_{i};d_{i})I_{n}f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}I_{n}f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}.$
Then by substituting (40) and (45) into (41), we can obtain the turbulent variance as
$\sigma_{tur,n}^{2}$ $\!=\!\frac{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}(\prod_{i=0}^{n}M_{2}(d_{i})-1)I_{n}f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}I_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}},$ (46)
where $M_{2}(d_{i})$ is the second-order moment of the turbulent fading coefficient $\eta_{i}(d_{i})$. For LN distribution, we have $M_{2}(d_{i})=\exp(\sigma_{r}^{2}(d_{i}))$; and for GG distribution, we have $M_{2}(d_{i})=(1+1/\alpha(d_{i}))(1+1/\beta(d_{i}))$.
Similar to the estimation of the average receiving power $P_{n}$, we can also use the MCI approach to estimate both the nominator and the denominator of $\sigma_{tur,n}^{2}$ in (46) using the same sampling PDF given in (25). Then using the same sampling points $\{\bm{s}_{n}^{1},\bm{s}_{n}^{2},\cdots,\bm{s}_{n}^{M}\}$, we can estimate the turbulent variance $\sigma_{tur,n}^{2}$ as
$\sigma_{tur,n}^{2}$ $=\frac{\frac{1}{M}\sum_{m=1}^{M}\left[\prod_{i=0}^{n}M_{2}(d_{i}^{m})-1\right]I_{n}(\bm{s}_{n}^{m})}{\frac{1}{M}\sum_{m=0}^{M}I_{n}(\bm{s}_{n}^{m})}=\frac{1}{Count}\sum_{m=1}^{M}\left[\prod_{i=0}^{n}M_{2}(d_{i}^{m})-1\right]I_{n}(\bm{s}_{n}^{m}),$ (47)
where $Count\triangleq\sum_{m=0}^{M}I_{n}(\bm{s}_{n}^{m})$ is the number of sampling points satisfying $\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F}$.
Similar to the $n$th scattering order, the total turbulent fluctuation effect on the total receiving power of $N$ scattering orders can be characterized by defining an equivalent turbulent fading coefficient $\eta_{eq}\triangleq\sum_{n=1}^{N}P_{tur,n}(\bm{\eta})/\sum_{n=1}^{N}P_{n}$. Using the relations $P_{tur,n}(\bm{\eta})=\eta_{eq,n}P_{n}$ and $P_{tot}=\sum_{n=1}^{N}P_{n}$, we can obtain
$\eta_{eq}$ $=\sum_{n=1}^{N}\frac{P_{n}}{P_{tot}}\eta_{eq,n}.$ (48)
Then the total turbulent variance $\sigma_{tur}^{2}$ for $N$ scattering orders can be obtained as
$\sigma_{tur}^{2}$ $\triangleq\text{Var}[\eta_{eq}]=\sum_{n=1}^{N}\left(\frac{P_{n}}{P_{tot}}\right)^{2}\sigma_{tur,n}^{2},$ (49)
where $P_{n}$, $P_{tot}$, and $\sigma_{tur,n}^{2}$ are given by (32), (33), and (47), respectively.
IV-D Estimating the Distribution of Turbulent Fading Coefficient
It is challenging to derive the distribution of the turbulent fading coefficient for a general turbulent fading model. However, when the turbulence is weak and an LN fading is considered, then we can further derive the distributions of the equivalent turbulent fading coefficients $\eta_{eq,n}$ and $\eta_{eq}$. For a given propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$, the equivalent turbulent fading coefficient $\eta_{eq,n}$ can be obtained from (40) as
$\eta_{eq,n}(\bm{d},\bm{\theta},\bm{\phi})=\prod_{i=0}^{n}\eta_{i}(\bm{d},\bm{\theta},\bm{\phi}).$ (50)
The logarithm of $\eta_{eq,n}(\bm{d},\bm{\theta},\bm{\phi})$ is
$\ln\eta_{eq,n}(\bm{d},\bm{\theta},\bm{\phi})=\sum_{i=0}^{n}\ln\eta_{i}(\bm{d},\bm{\theta},\bm{\phi}).$ (51)
Because $\eta_{i}(\bm{d},\bm{\theta},\bm{\phi})$ satisfies an LN distribution with parameter $\sigma_{I}^{2}(d_{i})$ and $\{\eta_{0},\eta_{1},\cdots,\eta_{n}$ are independent variables, from (51) we can see that $\eta_{eq,n}(\bm{d},\bm{\theta},\bm{\phi})$ also satisfies an LN distribution with parameter $\sigma_{eq,n}^{2}(\bm{d},\bm{\theta},\bm{\phi})=\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i})$. Therefore, the conditional PDF of $\eta_{eq,n}(\bm{d},\bm{\theta},\bm{\phi})$ is given by
$f_{H,n}(\eta_{eq,n}|\bm{d},\bm{\theta},\bm{\phi})$ $=\frac{1}{\eta_{eq,n}\sqrt{2\pi\sigma_{eq,n}^{2}(\bm{d},\bm{\theta},\bm{\phi})}}\exp\left(\!{-\frac{(\ln\eta_{eq,n}+0.5\sigma_{eq,n}^{2}(\bm{d},\bm{\theta},\bm{\phi}))^{2}}{2\sigma_{eq,n}^{2}(\bm{d},\bm{\theta},\bm{\phi})}}\!\right).$ (52)
Then the unconditional PDF for the equivalent turbulent fading coefficient $\eta_{eq,n}$ can be obtained by averaging $\{\bm{d},\bm{\theta},\bm{\phi}\}$ on $\mathcal{F}$ out, which is given by
$f_{H,n}(\eta_{eq,n})$ $=\frac{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}\frac{I_{n}(\bm{d},\bm{\theta},\bm{\phi})}{\eta_{eq,n}\sqrt{2\pi\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i})}}\exp\left({-\frac{(\ln\eta_{eq,n}+0.5\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i}))^{2}}{2\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i})}}\right)f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}I_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}.$ (53)
Then we can also use the MCI approach to estimate $f_{H}(\eta_{eq,n})$ using the same sampling PDF given in (25). Using the same sampling points $\{\bm{s}_{n}^{1},\bm{s}_{n}^{2},\cdots,\bm{s}_{n}^{M}\}$, we can estimate the PDF of equivalent turbulent fading coefficient $f_{H}(\eta_{eq,n})$ as
$f_{H,n}(\eta_{eq,n})$ $=\frac{1}{Count}\sum_{m=1}^{M}\frac{I_{n}(\bm{s}_{n}^{m})}{\eta_{eq,n}\sqrt{2\pi\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i}^{m})}}\exp\left({-\frac{(\ln\eta_{eq,n}+0.5\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i}^{m}))^{2}}{2\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i}^{m})}}\right).$ (54)
According to (48), the PDF of the total equivalent turbulent fading coefficient $\eta_{eq}$ for $N$ scattering orders can be obtained as the convolution of the PDFs of $\{\frac{P_{1}}{P_{tot}}\eta_{eq,1},\frac{P_{2}}{P_{tot}}\eta_{eq,2},\cdots,\frac{P_{N}}{P_{tot}}\eta_{eq,N}\}$, i.e.,
$f_{H}(\eta_{eq})$ $=\frac{P_{tot}}{P_{1}}f_{H,1}\left(\frac{P_{tot}}{P_{1}}\eta_{eq}\right)*\frac{P_{tot}}{P_{2}}f_{H,2}\left(\frac{P_{tot}}{P_{2}}\eta_{eq}\right)*\cdots*\frac{P_{tot}}{P_{N}}f_{H,N}\left(\frac{P_{tot}}{P_{N}}\eta_{eq}\right),$ (55)
where * denotes the convolution operation and $f_{H,n}(x)$ is calculated by (54).
The pseudocode of the MCI approach is similar to the one we presented in [17]. The only difference is that we need to add the estimation of turbulent fluctuation and distribution in each MCI process according to the equations given in (47) and (54), respectively.
V Numerical Results and Discussions
In this section, we present some numerical results to verify our theoretical analysis and discuss some interesting findings. Unless otherwise specified, the simulation parameters we used in this section are listed in Table I. Without loss of generality, we set $C_{n}^{2}=10^{-17}$ $\text{m}^{-2/3}$, $C_{n}^{2}=10^{-15}$ $\text{m}^{-2/3}$, and $C_{n}^{2}=10^{-13}$ $\text{m}^{-2/3}$ for weak, moderate, and strong turbulent conditions, respectively.
TABLE I: Simulation Parameters
Parameters Values
$r$ $500$ m
$[\theta_{T},\phi_{T}]$ $[45^{\circ},-90^{\circ}]$
$[\theta_{R},\phi_{R}]$ $[45^{\circ},90^{\circ}]$
$[\beta_{T},\beta_{R}]$ $[17^{\circ},30^{\circ}]$
$A_{r}$ $1.77\times 10^{-4}{\text{m}}^{2}$
$\lambda$ $2.6\times 10^{-7}$ m
[…]
$[k_{a}^{par},k_{s}^{ray},k_{s}^{mie}]$ $[0.802,0.266,0.284]$ ${\text{km}}^{-1}$
$[\gamma,g,f]$ $[0.017,0.72,0.5]$
$L_{0}$ $10^{2}$ m
$d_{0}$ $10^{-3}$ m
$C_{n}^{2}$ $10^{-15}$ $\text{m}^{-2/3}$
$\Delta t$ $0.02$ $\mu$s
$M$ $10^{6}$
$N$ $3$
V-A Results for Turbulence-Induced Scattering Effect
Figure 2: Turbulence-induced scattering effect: (a) scattering coefficient $k_{s}^{tur}$ under different $C_{n}^{2}$ and $d_{0}$; (b) phase scattering function $p^{tur}(\theta_{s})$ at different $\theta_{s}$ and $d_{0}$
We first explore the turbulence-induced scattering coefficient $k_{s}^{tur}$ and phase scattering function $p^{tur}(\theta_{s})$ under different turbulent parameters in Fig. 2. Fig. 2(a) presents the scattering coefficient $k_{s}^{tur}$ under different refractive-index structure parameter $C_{n}^{2}$ and correlation distance $d_{0}$. From Fig. 2(a) we can see that the scattering coefficient $k_{s}^{tur}$ increases as either the turbulence strength or the correlation distance increases. Besides, for a given refractive-index structure parameter $C_{n}^{2}$, the slop of the turbulent scattering coefficient $k_{s}^{tur}$ for $d_{0}\gg\lambda$ is much smaller than that for $d_{0}\ll\lambda$, where $\lambda=2.6\times 10^{-7}$ m. These observations verified our theoretical analysis on Eq. (6). Fig. 2(b) presents the phase scattering function $p^{tur}(\theta_{s})$ at different scattering angle $\theta_{s}$ and correlation distance $d_{0}$. Note that $L_{0}$ and $C_{n}^{2}$ have no impact on the phase scattering function. From Fig. 2(b), we can see that, when $d_{0}\gg\lambda$, the scattering intensity rapidly increases as the scattering angle decreases. This indicates that the incident photons will have a high probability of maintaining their original traveling direction after scattering. By contrast, when $d_{0}\ll\lambda$, we can see that the phase scattering function varies slowly at different scattering angles and will finally approximate an isotropic scattering. These observations verified our theoretical analysis on Eq. (7).
The impacts of turbulence-induced scattering effect on the receiving power of NLOS UV communications should be explored by considering both the scattering coefficient $k_{s}^{tur}$ and the phase scattering function $p^{tur}(\theta_{s})$. Combining the results in Figs. 2(a) and 2(b), we have the following observations:
  • When $d_{0}\ll\lambda$, we have $k_{s}^{tur}\to 0$; in this case, though the phase scattering function approaches an isotropic scattering, the overall impact of the turbulence-induced scattering effect is negligible compared with the scattering effect due to particles; therefore, we can always ignore the turbulence-induced scattering effect when $d_{0}\ll\lambda$.
  • When $d_{0}\gg\lambda$, the scattering is almost a complete forward scattering since almost all the scattering intensity concentrates at $\theta_{s}=0^{\circ}$; in this case, though there exists large scattering coefficient $k_{s}^{tur}\gg 1$, the overall turbulence-induced scattering can be equivalently regarded as a rectilinear propagation with negligible divergence angle, which means no evident impact on the final receiving power can be observed; therefore, we can also ignore the turbulence-induced scattering effect when $d_{0}\gg\lambda$.
  • When $d_{0}\sim\lambda$, the phase scattering function is not a complete forward scattering; then we cannot ignore the turbulence-induced scattering effect if the scattering coefficient $k_{s}^{tur}$ is not negligible compared with $k_{s}^{par}$, which can happen under strong turbulent conditions; for example, when $d_{0}=10^{-7}$ m, $C_{n}^{2}=10^{-13}$ $\text{m}^{-2/3}$, and $L_{0}=10^{3}$ m, we have $k_{s}^{tur}=0.586$ $\text{km}^{-1}$, which is comparable with $k_{s}^{par}=0.550$ $\text{km}^{-1}$.
Figure 3: CIR results under different outer scale parameters $L_{0}$: (a) $r=200$ m; (b) $r=500$ m; (c) $r=800$ m
Let us explore the cases with $d_{0}\sim\lambda$. Fig. 3 presents the CIR results under different outer scale parameters $L_{0}$ in different communication distances $r\in\{200\ \text{m},500\ \text{m},800\ \text{m}\}$ with $d_{0}=10^{-7}$ m and $C_{n}^{2}=10^{-13}$ $\text{m}^{-2/3}$. Comparing Figs. 3(a), 3(b), and 3(c), we can see that when the communication distance is small, the turbulence-induced scattering can enhance the receiving power. However, as the communication distance increases, the enhancement due to the turbulence-induced scattering gradually decreases and finally the turbulence-induced scattering will introduce an extra attenuation on receiving power in large communication distance. Here we can explain this distance-dependent scaling effect on the receiving power. Since the single-scattering is the major scattering here, we roughly consider the single-scattering only. As we know, an increase in the scattering coefficient can not only increase the scattering strength but also increase the extinction strength and shorten the photon traveling distance. The expectation of the photon traveling distance can be estimated as $1/k_{e}^{tot}$ and here we have $1/k_{e}^{tot}=1/(k_{s}^{par}+k_{a}^{par})\approx 739.6$ m by ignoring the turbulence-induced scattering. Then when the communication distance is smaller than $739.6$ m, an increase in the scattering coefficient will shorten the expected traveling distance and thus increase the scattering strength in the common volume, and vice versa. Therefore, when $r=200$ m, the turbulence-induced scattering mainly acts as a scattering effect enhancer; whereas when $r=800$ m, the turbulence-induced scattering mainly acts as an extinction effect enhancer.
Though we have demonstrated that the turbulence-induced scattering effect can impact the final receiving power when $d_{0}\sim\lambda$, in practical implementations, we have $d_{0}\gg\lambda$ almost always holds. For example, the correlation distance $d_{0}$ is normally in between $l_{0}$ and $L_{0}$, where $l_{0}$ is the inner scale parameter representing the smallest turbulent eddy size; and $l_{0}$ is at the same order of magnitude with the Fresnel zone $R_{F}=(L/k)^{1/2}$, where $k$ is the light wave number and $L$ is the light propagating distance. For typical UV communication scenarios, we have $10^{0}\ \text{m}\leq L\leq 10^{4}\ \text{m}$ and $2.034\times 10^{-4}\ \text{m}\leq R_{F}\leq 2.034\times 10^{-2}\ \text{m}$, which results in $d_{0}\sim R_{F}\gg\lambda$. Therefore, we can conclude that the turbulence-induced scattering effect can always be ignored in typical UV communication scenarios.
V-B Results for Turbulent Fluctuation Effect
Figure 4: Turbulent fluctuations under different number of sampling points
Then we explore the turbulent fluctuation effects. Fig. 4 presents the estimated turbulent fluctuation under different number of sampling points $M$. As we can see from Fig. 4, the estimated turbulent fluctuation $\sigma^{2}_{tur}$ or $\sigma_{tur,n}^{2}$ will become stable as $M$ increases. Besides, we can also roughly estimate the turbulent variance by using the method proposed for the single-scattering link in [47]. Specifically, the turbulent variance can be approximated as $\sigma_{tur,1}^{2}\approx\exp(\sigma_{I}^{2}(r_{1})+\sigma_{I}^{2}(r_{2}))-1\approx 0.0493$, which is quite close to our result $\sigma_{tur,1}^{2}=0.0492$ in Fig. 4. This verified that our proposed estimation method for turbulent fluctuation can well estimate the turbulent variance.
Figure 5: Turbulent fluctuations under various system geometries: (a) turbulent fluctuations under different communication distances; (b) turbulent fluctuations under different zenith angles; (c) turbulent fluctuations under different divergence angles
Then we present the turbulent fluctuations results under various system geometries in Fig. 5. From Figs. 5(a), 5(b) and 5(c), we can see that the turbulent variance increases as either the communication distance $r$, the elevation angle $\pi/2-\theta_{T}$, or the divergence angle $\beta_{T}$ increases, which is compatible with reported experimental results in [53, 54]. Besides, we can also observe that the turbulence variance will approach a stable value for strong turbulent conditions in Figs. 5(a), 5(b) and 5(c). This is because the channel will become a saturated turbulent channel under strong turbulence as either the communication distance $r$, the elevation angle $\pi/2-\theta_{T}$, or the divergence angle $\beta_{T}$ increases, which is compatible with the saturated turbulence case with $\sigma_{r}^{2}\to\infty$ at the end of Section II.
Figure 6: Distribution of turbulent fading coefficient under different distances: (a) $r=200$ m; (b) $r=500$ m; (c) $r=800$ m
At last, we present the estimated probability density for the turbulent fading coefficients under different communication distances $r\in\{200\ \text{m},500\ \text{m},800\ \text{m}\}$ with $C_{n}^{2}=10^{-17}$ $\text{m}^{-2/3}$ in Fig. 6. The simulation results of $f_{H,n}(\eta_{eq,n})$ for the $n$th order scattering are fitted with a weighted summation of $n$ Gaussian distributions, and the simulation results of $f_{H}(\eta_{eq})$ are fitted with a Gaussian distribution. From Fig. 6, we can see that the Gaussian fitting can well match the simulation results. Here we give an explanation on the single-scattering only. For a single-scattering, we can divide the common volume $V$ into many small volumes $\{v_{1},v_{2},\cdots,v_{L}\}$. Then according to the LN assumption in [47], the receiving power $P_{r}^{l}$ contributed from the $l$th volume $v_{l}$ follows a LN distribution. Then according to the central limit theorem, the total receiving power $P_{r}=\sum_{l=1}^{L}P_{r}^{l}$ will converge to a Gaussian distribution when $L$ is large.
VI Conclusion
Existing works on the modeling of multiple-scattering turbulent channels either ignored the turbulence-induced scattering effect or erroneously estimated the turbulent fluctuation effect on the NLOS UV communications. In this paper, we established a multiple-scattering turbulent channel model considering both the turbulence-induced scattering effect and the turbulent fluctuation effect. Numerical results demonstrated that, for typical UV communication scenarios, the turbulence-induced scattering effect can always be ignored; and the turbulent fluctuation will increase as either the communication distance, the elevation angle, or the divergence angle increases, which is compatible with existing experimental results. Besides, we also demonstrated that the distribution of the overall turbulent fading coefficient for the multiple-scattering process can be approximated as a Gaussian distribution.

### v3

@@
-Monte-Carlo Integration Based Multiple-Scattering Channel Modeling for Ultraviolet Communications in Turbulent Atmosphere
+Turbulent Multiple-Scattering Channel Modeling for Ultraviolet Communications: A Monte-Carlo Integration Approach
 Abstract
 Modeling of multiple-scattering channels in atmospheric turbulence is essential for the performance analysis of long-distance non-line-of-sight (NLOS) ultraviolet (UV) communications.
-Existing works on the turbulent channel modeling for NLOS UV communications either ignored the turbulence-induced scattering effect or erroneously estimated the turbulent fluctuation effect, resulting in a contradiction with reported experiments.
-In this paper, we establish a comprehensive multiple-scattering turbulent channel model for NLOS UV communications considering both the turbulence-induced scattering effect and the turbulent fluctuation effect.
-We first derive the turbulent scattering coefficient and turbulent phase scattering function based on the Booker-Gordon turbulent power spectral density model.
-Then an improved estimation method is proposed for both the turbulent fluctuation and the turbulent fading coefficient based on the Monte-Carlo integration approach.
+Existing works on the turbulent channel modeling for NLOS UV communications either focused on single-scattering cases or estimate the turbulent fluctuation effect in an unreliable way based on Monte-Carlo simulation (MCS) approach.
+In this paper, we establish a comprehensive turbulent multiple-scattering channel model by using a more efficient Monte-Carlo integration (MCI) approach for NLOS UV communications, where both the scattering, absorption, and turbulence effects are considered.
+Compared with the MCS approach, the MCI approach is more interpretable for estimating the turbulent fluctuation.
+To achieve this, we first introduce the scattering, absorption, and turbulence effects for NLOS UV communications in turbulent channels.
+Then we propose the estimation methods based on MCI approach for estimating both the turbulent fluctuation and the distribution of turbulent fading coefficient.
 Numerical results demonstrate that the turbulence-induced scattering effect can always be ignored for typical UV communication scenarios.
-Besides, the turbulent fluctuation will increase as either the communication distance, the elevation angle, or the divergence angle increases, which is compatible with existing experimental results.
-Moreover, we find that the probability density of the equivalent turbulent fading for multiple-scattering turbulent channels can be approximated as a Gaussian distribution.
+Besides, the turbulent fluctuation will increase as either the communication distance increases or the zenith angle decreases, which is compatible with existing experimental results and also with our experimental results.
+Moreover, we demonstrate numerically that the distribution of the turbulent fading coefficient for UV multiple-scattering channels under all turbulent conditions can be approximated as log-normal distribution; and we also demonstrate both numerically and experimentally that the turbulent fading can be approximated as a Gaussian distribution under weak turbulence.
 Index Terms:
@@
 I-A Background and Motivation
-The ultraviolet (UV) communication [1, 2] employing the “solar-blind” UV signals ($200$-$280$ nm) as the information carriers has attracted increasing attention in recent decades.
-Due to its inherent advantages including low background noise, high local security, good adaptivity to extreme weather, and ability of non-line-of-sight (NLOS) links, the UV communication becomes a promising technique for communication scenarios where radio-frequency (RF) communications are impermissible or line-of-sight (LOS) links for traditional free-space optical (FSO) are unavailable [1, 2].
-Considerable works have been performed on the NLOS UV channel modeling [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], channel estimation [21, 22, 23, 24, 25, 26], coding and modulation [27, 28, 29, 30], diversity reception [31, 32, 33], duplex and relay techniques [34, 35, 36, 37], and experimental tests [38, 39, 40, 41, 42, 43].
-However, most existing works [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] on the channel modeling ignored the impacts of atmospheric turbulence, which cannot be applied to the turbulent cases under either long communication distances or large elevation angles.
-The major effects of the atmospheric turbulence on the photon propagating process can be divided into two types: the turbulence-induced scattering effect when a photon interacts with the atmospheric medium [44] and the turbulent fluctuation effect when a photon travels in the atmosphere [45].
-In existing works [46, 47, 48, 49, 50, 51, 52], the turbulence-induced scattering effects [46] and the turbulent fluctuation effects [47, 48, 49, 50, 51, 52] were separately studied in the channel modeling of NLOS UV communications.
-Besides, there exist some contradictions between the existing theoretical turbulent fluctuation model [52] and experimental results [53, 54].
-Therefore, a comprehensive turbulent channel model considering both the turbulence-induced scattering effect and the turbulent fluctuation effect is urgently demanded for long-distance NLOS UV communications.
+The ultraviolet (UV) communication [yuan2016review] employing the “solar-blind” UV signals (200200-$280$ nm) as the information carriers has attracted increasing attention in recent decades.
+Due to its inherent advantages including low background noise, high local security, good adaptivity to extreme weather, and ability of non-line-of-sight (NLOS) links, the UV communication becomes a promising technique for communication scenarios where radio-frequency (RF) communications are impermissible or line-of-sight (LOS) links for traditional free-space optical (FSO) are unavailable [yuan2016review].
+In recent decade, considerable works have been performed on the NLOS UV channel modeling [yuan2016integral, wu2019single, shen2019modeling, cao2021single, cao2022single, ding2009modeling, drost2011uv, ding2010path, yuan2019importance, yuan2019monte, shan2020modeling], channel estimation [gong2015channel, gong2016optical, wei2018simultaneous, hu2020non, shen2021lmmse], coding and modulation [cao2023performance, cao2023power], diversity reception [qin2017noncoplanar, yuan2023simo, wang2023mimo], duplex and relay techniques [ardakani2017performance, gong2018full, wang2023nlos], and experimental tests [sun201771, wang20181mbps, alkhazragi2020gbit].
+However, most existing works on the channel modeling ignored the impacts of atmospheric turbulence, which cannot be applied to the turbulent cases under either long communication distances or large elevation angles.
+In existing literatures, the major effects of the atmospheric turbulence on the photon propagating process can be divided into two types: the turbulence-induced scattering effect when a photon interacts with the atmospheric medium [ishimaru2005theory] and the turbulent fluctuation effect when a photon travels in the atmosphere [andrews2005laser].
+In existing works, the turbulence-induced scattering effects and the turbulent fluctuation effects were separately studied in the channel modeling of NLOS UV communications.
+Besides, existing works on turbulent effects in NLOS UV communications either focused on single-scattering cases or estimate the turbulent fluctuation based on a Monte-Carlo simulation (MCS) approach.
+As it is demonstrated in [yuan2019importance, yuan2019monte], the Monte-Carlo integration (MCI) approach can provide a more efficient modeling for multiple scattering channels.
+Besides, as we will demonstrate later in this paper, the MCI approach can also provide a more interpretable logic for estimating the turbulent fluctuation.
+Therefore, a comprehensive turbulent multiple scattering channel modeling based on MCI approach considering both the scattering, absorption, and turbulence effects is urgently demanded for long-distance NLOS UV communications.
 I-B Related Works
-The turbulence-induced scattering effect on light wave propagation in the atmosphere was elaborately studied and summarized in [44], where the turbulent atmosphere is modeled as a random continuum with refractive-index varying randomly and continuously in time and space.
-Some preliminary results on the receiving power and channel impulse response (CIR) of NLOS optical links were derived in [44] based on a single-scattering assumption.
-The impact of the turbulence-induced scattering effect on the NLOS UV communication was first studied in [46], where the authors claimed that more photons are scattered to the receiver in stronger turbulence for single-scattering cases.
+The turbulence-induced scattering effect on light wave propagation in the atmosphere was elaborately studied and summarized in [ishimaru2005theory], where the turbulent atmosphere is modeled as a random continuum with refractive-index varying randomly and continuously in time and space.
+Some preliminary results on the receiving power and channel impulse response (CIR) of NLOS optical links were derived in [ishimaru2005theory] based on a single-scattering assumption.
+The impact of the turbulence-induced scattering effect on the NLOS UV communication was first studied in [xiao2013non].
 However, the impact of the turbulence-induced scattering on the phase scattering function for NLOS UV communications was not clearly studied.
-Besides, a recent study [55] in underwater optical communication also incorporated the turbulence-induced scattering effect into the scattering-based channel model and demonstrated an addition of more than $5$ dB path loss when the turbulence is considered, which seems opposite to the numerical results obtained in [46].
+Besides, a recent study [xu2022improvement] in underwater optical communication also incorporated the turbulence-induced scattering effect into the scattering-based channel model and demonstrated an addition of more than $5$ dB path loss when the turbulence is considered, which seems opposite to the numerical results obtained in [xiao2013non].
 Therefore, the impact of the turbulence-induced scattering effect on the NLOS UV communication is still not clearly indicated.
-The turbulent fluctuation effect on light wave propagating for LOS optical links was elaborately studied and summarized in [45].
-The impact of turbulent fluctuation effect on the NLOS UV communications was first studied in a single-scattering model under small common volume assumption [47], where the NLOS link was divided into two LOS links: one from the transmitter to the common volume and another from the common volume to the receiver.
-When two turbulent fluctuations with log-normal (LN) distribution are introduced into these two LOS links, the authors in [47] demonstrated that the overall turbulent fluctuation can also be approximated as an LN distribution.
-Based on [47], the bit-error rate performance of NLOS UV communications using this turbulent fluctuation model was studied in [48].
-Besides, the turbulent single-scattering model in [47] was extended to narrow beam cases in [49].
-A turbulence-induced attenuation effect for LOS links was introduced into the turbulent channel model of NLOS UV communications in [49, 50, 51].
-However, the turbulence-induced attenuation considered in LOS laser communications comes from the beam wandering and beam spreading effects [56] caused by the misalignment between the transceivers.
-Therefore, turbulence-induced attenuation does not apply to NLOS UV links since no alignment is required in NLOS UV communications.
-Besides, these works [47, 48, 49, 50, 51] are based on the single-scattering model, which cannot be applied to multiple-scattering based long-distance NLOS UV communications.
-The turbulent fluctuation model for multiple-scattering was first studied in [52], where the Monte-Carlo based multiple-scattering model [13, 14, 15, 16, 17, 18, 19] is adopted and a Gamma-Gamma (GG) turbulent fading is introduced into each LOS link between two scatters.
-However, the authors in [52] estimated the turbulent fluctuation by simply calculating the variance of the average receiving power, which cannot really capture the turbulent fluctuation effect.
-This is because the randomness of the obtained average receiving power in a Monte-Carlo based channel model comes from not only the turbulent fluctuation but also the randomness of the Monte-Carlo process itself [15, 17].
-As we will demonstrate in Section IV-C, any attempt to calculate the turbulent fluctuation by estimating the average receiving power of Monte-Carlo based channel models will fail.
-Based on the same reason, the distribution of the receiving power obtained in [52] cannot characterize the turbulent fading either.
-Besides, simulation results on the turbulent fluctuation obtained in [52] indicated that the multiple-scattering effect can somehow mitigate the turbulent fluctuation effect, which is opposite to reported experimental results [53, 54].
-Therefore, a more sophisticated turbulent fluctuation model is demanded for characterizing the multiple-scattering turbulent channel.
+The turbulent fluctuation effect on light wave propagating for LOS optical links was elaborately studied and summarized in [andrews2005laser].
+The impact of turbulent fluctuation effect on the NLOS UV communications was first studied in a single-scattering model under small common volume assumption [ding2011turbulence], where the NLOS link was divided into two LOS links: one from the transmitter to the common volume and another from the common volume to the receiver.
+When two turbulent fluctuations with log-normal (LN) distribution are introduced into these two LOS links, the authors in [ding2011turbulence] demonstrated that the overall turbulent fluctuation can also be approximated as an LN distribution.
+Based on [ding2011turbulence], the bit-error rate performance of NLOS UV communications using this turbulent fluctuation model was studied in [liu2015performance].
+Besides, the turbulent single-scattering model in [ding2011turbulence] was extended to narrow beam cases in [shan2020single].
+A turbulence-induced attenuation effect for LOS links was introduced into the turbulent channel model of NLOS UV communications in [shan2020single, zuo2012effect, xiao2012non].
+However, the turbulence-induced attenuation considered in LOS laser communications comes from the beam wandering and beam spreading effects [kaushal2017free] caused by the misalignment between the transceivers.
+Therefore, turbulence-induced attenuation cannot directly be applied to NLOS UV links since no alignment is required in NLOS UV communications.
+Recently, a turbulent channel modeling based on equivalent scattering points approach was proposed in [chu2024turbulent], where a Gaussian distribution on the turbulent fading coefficient was obtained.
+However, these works [ding2011turbulence, liu2015performance, shan2020single, zuo2012effect, xiao2012non, chu2024turbulent] are based on the single-scattering model, which cannot be applied to multiple-scattering based long-distance NLOS UV communications.
+The turbulent fluctuation model for multiple-scattering was first studied in [wang2013characteristics] based on an MCS approach [drost2011uv], where a Gamma-Gamma (GG) turbulent fading is introduced into each LOS link between two scatters.
+However, the authors in [wang2013characteristics] estimated the turbulent fluctuation by simply calculating the variance of the average receiving power.
+As we will demonstrate in Section IV-B, the attempt to calculate the turbulent fluctuation by estimating the average receiving power of Monte-Carlo based channel models is unreliable due to the properties of Monte-Carlo process.
+Besides, it was demonstrated that the MCI approaches enjoy more efficient modeling compared with the MCS approaches.
+Therefore, a more sophisticated turbulent fluctuation model based on the MCI approach is demanded for characterizing the turbulent multiple-scattering channel.
 I-C Contributions
-Our work aims to establish a multiple-scattering turbulent channel model considering both the turbulence-induced scattering effect and the turbulent fluctuation effect.
-To achieve this, we first derived the turbulent scattering coefficient and turbulent phase scattering function based on the Booker-Gordon turbulent power spectral density model.
-Then we proposed an improved estimation method based on a Monte-Carlo integration (MCI) approach for both the turbulent fluctuation and the distribution of the turbulent fading coefficient.
-The numerical results verified our theoretical analysis and some interesting results were obtained.
+Our work aims to establish a turbulent multiple-scattering channel model based on an MCI approach for NLOS UV communications, considering both scattering, absorption, and turbulence effects.
+To achieve this, we first introduce the scattering, absorption, and turbulence effects of NLOS UV communications in turbulent channels.
+Then we proposed an improved estimation method based on an MCI approach for both the turbulent fluctuation and the distribution of the turbulent fading coefficient.
+Both numerical simulation and experimental measurements are conducted to verified our theoretical analysis.
 Specifically, we summarize the main contributions of this work as follows:
-  • We established the first comprehensive multiple-scattering turbulent channel model considering both the turbulence-induced scattering effect and fluctuation effect for NLOS UV communications by using an MCI approach.
-  • We proposed an improved estimation method for turbulent fluctuation and further proposed an estimation method for the distribution of the turbulent fading coefficient for NLOS UV communications under weak turbulence.
-  • We demonstrated that the turbulence-induced scattering effect cannot be ignored only if the turbulent correlation distance approximates the light wavelength under strong turbulence.
-Besides, for typical UV communication scenarios, we demonstrated that the turbulence-induced scattering effect can always be ignored.
-  • We demonstrated that the turbulent fluctuation for NLOS UV communications will increase as either the communication distance, the elevation angle, or the divergence angle increases, which is compatible with existing experimental results.
-  • We demonstrated that the distribution of the turbulent fading coefficient for an $n$-order scattering under weak turbulence can be well fitted by a weighted summation of $n$ Gaussian distributions.
-Besides, the distribution of the overall turbulent fading coefficient for NLOS UV communications can be approximated as a Gaussian distribution.
+  • We established the first turbulent multiple-scattering channel model based on an MCI approach for NLOS UV communications, where the scattering, absorption, and turbulence effects are considered.
+  • We proposed an improved estimation method for turbulent fluctuation and further proposed an estimation method for the distribution of the turbulent fading coefficient for NLOS UV communications.
+  • We demonstrated that the turbulence-induced scattering effect can always be ignored for typical UV communication scenarios; and we demonstrated that the turbulent fluctuation for NLOS UV communications will increase as either the communication distance increases or the zenith angle decreases, which is compatible with existing experimental results and also with our experimental results.
+  • We demonstrated numerically that the distribution of the turbulent fading coefficient for NLOS UV communications in all turbulent conditions can be approximated by LN distributions; and we also demonstrated both numerically and experimentally that the turbulent fading under weak turbulence can be approximated as a Gaussian distribution.
 The rest of this paper is organized as follows.
@@
 Based on the photon propagating model, we then derive the receiving power of multiple-scattering turbulent channels in Section III.
-Then we introduce the estimation methods for both the average receiving power, the CIR, the turbulent fluctuation, and the distribution of the turbulent fading coefficient using an MCI approach in Section IV.
-We then present some numerical results in Section V and conclude our work in Section VI.
+Then we introduce the estimation methods based on an MCI approach for both the average receiving power, the turbulent fluctuation, and the distribution of the turbulent fading coefficient in Section IV.
+We then present some numerical and experimental results in Section V and conclude our work in Section VI.
 II Photon Propagating in Turbulent Channels
-The non-turbulent atmosphere can be regarded as a group of particles, which can be modeled as randomly distributed scatters including molecules and aerosols [44]; whereas the turbulent atmosphere is usually regarded as a group of turbulent eddies according to Kolmogorov’s theory of turbulence, which can be modeled as random continuum with refractive-index varying randomly and continuously in time and space [44].
-Similar to [55], we consider the scattering effects due to both the random scatterers and the random continuum in this paper.
+The non-turbulent atmosphere can be regarded as a group of particles, which can be modeled as randomly distributed scatters including molecules and aerosols [ishimaru2005theory]; whereas the turbulent atmosphere is usually regarded as a group of turbulent eddies according to Kolmogorov’s theory of turbulence, which can be modeled as random continuum with refractive-index varying randomly and continuously in time and space [ishimaru2005theory].
+Similar to [xu2022improvement], we consider the scattering effects due to both the random scatterers and the random continuum in this paper.
 Therefore, the potential scatters include both the particles and the turbulent eddies.
@@
 II-A1 Scattering Effect of Particles
-When the particle size is much smaller than the light wavelength, the scattering process of a light wave can be modeled as a Rayleigh scattering; whereas when the particle size is comparable to or larger than the light wavelength, the scattering process can be modeled as a Mie scattering [1, 2].
+When the particle size is much smaller than the light wavelength, the scattering process of a light wave can be modeled as a Rayleigh scattering; whereas when the particle size is comparable to or larger than the light wavelength, the scattering process can be modeled as a Mie scattering [yuan2016review].
 We denote the Rayleigh scattering coefficient and the Mie scattering coefficient by $k_{s}^{ray}$ and $k_{s}^{mie}$, respectively.
-Then the total scattering coefficient due to the randomly distributed particles is given by $k^{par}_{s}=k_{s}^{ray}+k_{s}^{mie}$.
-The scattering strength at different directions around the particle is characterized by the phase scattering function [44].
-For UV signals, the phase scattering functions for Rayleigh scattering and Mie scattering are respectively given by [1, 2] $p^{ray}(\theta_{s})=\frac{3[1+3\gamma+(1-\gamma)\cos^{2}\theta_{s}]}{16\pi(1+2\gamma)}$ and $p^{mie}(\theta_{s})=\frac{1-g^{2}}{4\pi}\left[\frac{1}{(1+g^{2}-2g\cos\theta_{s})^{{3}/{2}}}\right.$ $\left.+f\frac{(3\cos^{2}\theta_{s}-1)}{2(1+g^{2})^{{3}/{2}}}\right]$, where $\theta_{s}$ is the scattering (zenith) angle between the incident light direction and the scattering light direction; $\gamma$, $g$, and $f$ are the model parameters.
-The Rayleigh scattering approximates an isotropic scattering and the Mie scattering is a forward direction dominated scattering [1, 2].
+Then the total scattering coefficient due to the randomly distributed particles is given by
+$k^{par}_{s}=k_{s}^{ray}+k_{s}^{mie}.$ (1)
+The scattering strength at different directions around the particle is characterized by the phase scattering function [ishimaru2005theory].
+For UV signals, the phase scattering functions for Rayleigh scattering and Mie scattering are given by [yuan2016review]
+$\begin{cases}p^{ray}(\theta_{s})\!=\!\frac{3[1+3\gamma+(1-\gamma)\cos^{2}\theta_{s}]}{16\pi(1+2\gamma)},\\ p^{mie}(\theta_{s})\!=\!\frac{1-g^{2}}{4\pi}\!\!\left[\frac{1}{(1+g^{2}\!-\!2g\cos\theta_{s})^{{3}/{2}}}\!+\!\frac{f(3\cos^{2}\theta_{s}\!-\!1)}{2(1+g^{2})^{{3}/{2}}}\right],\end{cases}$ (2)
+where $\theta_{s}$ is the scattering (zenith) angle between the incident light direction and the scattering light direction; $\gamma$, $g$, and $f$ are the model parameters.
+The Rayleigh scattering approximates an isotropic scattering and the Mie scattering is a forward direction dominated scattering [yuan2016review].
 II-A2 Scattering Effect of Turbulence
-The turbulence-induced scattering effect for a light propagating in a random continuum can be characterized by the differential cross section per unit volume $\sigma(\bm{i},\bm{o})$ [44], where $\bm{i}$ and $\bm{o}$ are the unit vectors for the incident and the scattering directions, respectively.
-For a statistically homogeneous and isotropic random continuum and a scalar light wave, the differential cross section $\sigma(\bm{i},\bm{o})$ can be obtained as [44]
-$\sigma(\bm{i},\bm{o})=2\pi k^{4}\Phi_{n}(k_{s}),$ (1)
-where $\Phi_{n}(k_{s})$ is the power spectral density of the turbulence; $k={2\pi}/{\lambda}$ is the wave number of the incident light; $k_{s}=2k\sin({\theta_{s}}/{2})$ and $\theta_{s}$ is the scattering angle between $\bm{i}$ and $\bm{o}$.
-Then the scattering coefficient caused by the turbulence can be obtained as [44]
-$k_{s}^{tur}=\int_{\Omega}\sigma(\bm{i},\bm{o})\mathrm{d}\omega=4\pi^{2}k^{2}\int_{0}^{2k}\Phi_{n}(k_{s})k_{s}\mathrm{d}k_{s},$ (2)
-where $\mathrm{d}\omega$ is the differential solid angle and $\Omega$ is the whole solid angle.
-From (1) we can observe that the differential cross section $\sigma(\bm{i},\bm{o})$ is a function of the relative direction between $\bm{i}$ and $\bm{o}$.
-Then we can rewrite $\sigma(\bm{i},\bm{o})$ as $\sigma(\theta_{s})=2\pi k^{4}\Phi_{n}\left(2k\sin({\theta_{s}}/{2})\right)$ and further define the phase scattering function due to the turbulence as
-$p^{tur}(\theta_{s})$ $\triangleq\frac{\sigma(\theta_{s})}{\int_{\Omega}\sigma(\theta_{s})\mathrm{d}\omega}=\frac{2\pi k^{4}}{k_{s}^{tur}}\Phi_{n}\left(2k\sin({\theta_{s}}/{2})\right).$ (3)
-The scattering coefficient $k_{s}^{tur}$ in (2) and the phase scattering function $p^{tur}(\theta_{s})$ in (3) are closely related to the power spectral density $\Phi_{n}(k_{s})$ of the turbulence.
-However, different assumptions on the turbulence can result in different forms of the power spectral density [44].
-Without loss of generality, we adopt the widely used Booker-Gordon model in this paper [44, 46].
-The Booker-Gordon model characterizes the turbulent medium by using two quantities, i.e., the variance of the refractive index $\langle n_{1}^{2}\rangle$ and the correlation distance $d_{0}$.
-Here $n_{1}\triangleq n-\langle n\rangle$ is the fluctuation of the refractive index $n$; $d_{0}$ represents the average size of the turbulent eddies.
-Then the correlation function of the refractive-index at two positions $\bm{r}_{1}$ and $\bm{r}_{2}$ in Booker-Gordon model is assumed to be [44, 46] $B_{n}(r_{d})\triangleq\langle n_{1}(\bm{r}_{1})n_{1}(\bm{r}_{2})\rangle=\langle n_{1}^{2}\rangle\exp\left({-{r_{d}}/{d_{0}}}\right)$, where $r_{d}=|\bm{r}_{1}-\bm{r}_{2}|$.
-According to the Wiener-Khinchin theorem [44, 46], the power spectral density $\Phi_{n}(k_{s})$ is the Fourier transform of the correlation function $B_{n}(r_{d})$.
-For a scalar wave, the power spectral density $\Phi_{n}(k_{s})$ can be obtained as [44, 46]
-$\Phi_{n}(k_{s})$ $=\frac{\langle n_{1}^{2}\rangle d_{0}^{3}}{\pi^{2}(1+k_{s}^{2}d_{0}^{2})^{2}}.$ (4)
-Then by substituting (4) into (2), we can derive the scattering coefficient of Booker-Gordon model as
-$k_{s}^{tur}$ $=\frac{8k^{4}\langle n_{1}^{2}\rangle d_{0}^{3}}{1+4k^{2}d_{0}^{2}}.$ (5)
-Besides, when the turbulence is assumed to be isotropic, we have [44, 46] $\langle n_{1}^{2}\rangle=\frac{C_{n}^{2}L_{0}^{{2}/{3}}}{1.91}$, where $L_{0}$ is the outer scale parameter representing the largest turbulent eddy size; $C_{n}^{2}$ is the refractive-index structure parameter representing the strength of the turbulence.
-The refractive-index structure parameter $C_{n}^{2}$ is defined as the constant coefficient of the refractive-index structure equation [44] $D_{n}(r)\triangleq\langle[n(\bm{r}_{1})-n(\bm{r}_{2})]^{2}\rangle=C_{n}^{2}r^{{2}/{3}}$, where $n(\bm{r}_{i})$ is the refractive-index at position $\bm{r}_{i}$ and $r=|\bm{r}_{1}-\bm{r}_{2}|$ is the distance between two positions $\bm{r}_{1}$ and $\bm{r}_{2}$.
-The refractive-index structure equation also provides a way for measuring $C_{n}^{2}$ in practical implementations.
-Typical values of $L_{0}$ can vary from the magnitude of $10^{0}$ m to $10^{3}$ m; typical values of $C_{n}^{2}$ are at the magnitude of $10^{-15}$ $\text{m}^{-2/3}$ and $C_{n}^{2}$ can vary from $10^{-17}$ for weak turbulence to $10^{-13}$ for strong turbulence [44, 46].
-Then by substituting $\langle n_{1}^{2}\rangle$ into (5), we can further obtain
-$k_{s}^{tur}=\frac{8k^{4}C_{n}^{2}L_{0}^{{2}/{3}}d_{0}^{3}}{1.91(1+4k^{2}d_{0}^{2})}.$ (6)
-From the expression of $k_{s}^{tur}$ in (6), we can observe that $k_{s}^{tur}\propto C_{n}^{2}$.
-Besides, when $k^{2}d_{0}^{2}\ll 1$, we have $k_{s}^{tur}\approx\frac{8k^{4}C_{n}^{2}L_{0}^{{2}/{3}}d_{0}^{3}}{1.91}$, which indicates that $k_{s}^{tur}\propto d_{0}^{3}$ when the correlation distance $d_{0}$ is much smaller than the wavelength $\lambda$.
-In this case, we also have $k_{s}^{tur}\propto 1/\lambda^{4}$, which coincides with the Rayleigh scattering.
-When $k^{2}d_{0}^{2}\gg 1$, we have $k_{s}^{tur}\approx\frac{2k^{2}C_{n}^{2}L_{0}^{{2}/{3}}d_{0}}{1.91}$, which indicates that $k_{s}^{tur}\propto d_{0}$ when the correlation distance $d_{0}$ is much larger than the wavelength $\lambda$.
-Substituting (4) and (5) into (3), we can obtain the phase scattering function for the Booker-Gordon model as
+The turbulence-induced scattering effect for a light propagating in a random continuum can be characterized by the differential cross section per unit volume $\sigma(\bm{i},\bm{o})$ [ishimaru2005theory], where $\bm{i}$ and $\bm{o}$ are the unit vectors for the incident and the scattering directions, respectively.
+For a statistically homogeneous and isotropic random continuum and a scalar light wave, the differential cross section $\sigma(\bm{i},\bm{o})$ can be obtained as $\sigma(\bm{i},\bm{o})=2\pi k^{4}\Phi_{n}(k_{s})$, where $\Phi_{n}(k_{s})$ is the power spectral density of the turbulence; $k={2\pi}/{\lambda}$ is the wave number of the incident light; $k_{s}=2k\sin({\theta_{s}}/{2})$ and $\theta_{s}$ is the scattering angle between $\bm{i}$ and $\bm{o}$.
+We can observe that the differential cross section $\sigma(\bm{i},\bm{o})$ is a function of the relative direction between $\bm{i}$ and $\bm{o}$.
+Then we can rewrite $\sigma(\bm{i},\bm{o})$ as $\sigma(\theta_{s})=2\pi k^{4}\Phi_{n}\left(2k\sin({\theta_{s}}/{2})\right)$ and further define the phase scattering function due to the turbulence as [ishimaru2005theory]
+$p^{tur}(\theta_{s})$ $\triangleq\frac{\sigma(\theta_{s})}{\int_{\Omega}\sigma(\theta_{s})\mathrm{d}\omega}=\frac{\Phi_{n}\left(2k\sin({\theta_{s}}/{2})\right)}{\int_{\theta_{s}}\Phi_{n}\left(2k\sin({\theta_{s}}/{2})\right)\mathrm{d}\theta_{s}}.$ (3)
+Without loss of generality, we adopt the widely used Booker-Gordon model [ishimaru2005theory, xiao2013non] with $\Phi_{n}(k_{s})=\frac{\langle n_{1}^{2}\rangle d_{0}^{3}}{\pi^{2}(1+k_{s}^{2}d_{0}^{2})^{2}}$, where $\langle n_{1}^{2}\rangle$ is the variance of refractive index; $d_{0}$ is the average size of turbulent eddies.
+When the turbulence is assumed to be isotropic, we have [ishimaru2005theory, xiao2013non] $\langle n_{1}^{2}\rangle=\frac{C_{n}^{2}L_{0}^{{2}/{3}}}{1.91}$, where $L_{0}$ is the outer scale parameter representing the largest turbulent eddy size; $C_{n}^{2}$ is the refractive-index structure parameter representing the strength of the turbulence.
+Then substituting $\Phi_{n}(k_{s})$ into (3), we can obtain the phase scattering function due to turbulence as
 $p^{tur}(\theta_{s})=\frac{1+4k^{2}d_{0}^{2}}{4\pi(1+4k^{2}d_{0}^{2}\sin^{2}({\theta_{s}}/{2}))^{2}}.$ (4)
-From the phase scattering function $p^{tur}(\theta_{s})$ in (7), we can observe that the phase scattering function depends only on the wave number $k$ and the correlation distance $d_{0}$.
-Besides, when $k^{2}d_{0}^{2}\ll 1$, we have $p^{tur}(\theta_{s})\approx 1/4\pi$, which indicates that the scattering will approach an isotropic scattering when the correlation distance $d_{0}$ is much smaller than the wavelength $\lambda$.
-This isotropic scattering property is similar to the Rayleigh scattering.
-When $k^{2}d_{0}^{2}\gg 1$, the majority of the scattering intensity happens at small scattering angles.
-This indicates that the scattering is forward direction dominated when the correlation distance $d_{0}$ is much larger than the wavelength $\lambda$, which is similar to the Mie scattering.
-Combining the coefficients $k_{s}^{ray}$, $k_{s}^{mie}$, and $k_{s}^{tur}$, we can obtain the total scattering coefficients as
-$k_{s}^{tot}=k_{s}^{ray}+k_{s}^{mie}+k_{s}^{tur}.$ (8)
-Similarly, combining the phase scattering functions $p^{ray}(\theta_{s})$, $p^{mie}(\theta_{s})$, and $p^{tur}(\theta_{s})$, we can obtain the total phase scattering function as a weighted function of $p^{ray}(\theta_{s})$, $p^{mie}(\theta_{s})$, and $p^{tur}(\theta_{s})$, i.e.,
-$p^{tot}(\theta_{s})=\frac{k_{s}^{ray}}{k_{s}^{tot}}p^{ray}(\theta_{s})+\frac{k_{s}^{mie}}{k_{s}^{tot}}p^{mie}(\theta_{s})+\frac{k_{s}^{tur}}{k_{s}^{tot}}p^{tur}(\theta_{s}).$ (9)
+Because for optical wavebands including UV wavebands, we always have $\lambda\ll d_{0}$, i.e., $k^{2}d_{0}^{2}\gg 1$.
+This indicates that the scattering happens always at forward direction, which is equivalent to a rectilinear propagation effect.
+Therefore, the turbulence induced scattering effect can be omitted since it has no true effect on the propagating direction of the light.
+At last, we can obtain the scattering coefficient in turbulent channels as
+$k_{s}^{tot}=k_{s}^{par}=k_{s}^{ray}+k_{s}^{mie}.$ (5)
+Similarly, combining the phase scattering functions $p^{ray}(\theta_{s})$ and $p^{mie}(\theta_{s})$, we can obtain the total phase scattering function in turbulent channel as
+$p^{tot}(\theta_{s})=\frac{k_{s}^{ray}}{k_{s}^{tot}}p^{ray}(\theta_{s})+\frac{k_{s}^{mie}}{k_{s}^{tot}}p^{mie}(\theta_{s}).$ (6)
 II-B Absorption Effects in Turbulent Channels
-[…]
+Besides the scattering effects, the light wave can also be selectively absorbed by the particles due to the electronic transition between different energy levels.
+The absorption coefficient caused by the turbulence depends on the complex dielectric constant $\epsilon=\epsilon_{r}+i\epsilon_{i}$, i.e., $k_{a}^{tur}=k\epsilon_{i}\epsilon_{r}^{-\frac{1}{2}}$, where $\epsilon_{i}=60\lambda\delta$ and $\delta$ is the atmospheric conductivity.
+For the atmosphere near the ground, we have $\epsilon_{r}\approx 1.00059$ and $\delta\approx 2.2\times 10^{-14}$ $\text{S}/\text{m}$.
+Then we can obtain the absorption coefficient due to the turbulence as $k_{a}^{tur}\approx 4.1457\times 10^{-12}$, which is much smaller than the absorption coefficient due to the particles.
+Therefore, we can always ignore the absorption effect due to the turbulence for UV communications.
 Denoting the absorption coefficient due to the particles by $k_{a}^{par}$, we can obtain the total extinction coefficient of the atmosphere as
-$k^{tot}_{e}\triangleq k_{s}^{tot}+k^{par}_{a}=k_{s}^{ray}+k_{s}^{mie}+k_{s}^{tur}+k^{par}_{a}.$ (10)
+$k^{tot}_{e}\triangleq k_{s}^{tot}+k^{par}_{a}=k_{s}^{ray}+k_{s}^{mie}+k^{par}_{a}.$ (7)
 II-C Fluctuation Effects in Turbulent Channels
-Besides the scattering and absorption effects, the turbulence also introduces a fluctuation effect on the light irradiance [45].
-The turbulent fluctuation effect can be characterized by the turbulent fading coefficient $\eta$ with $\eta\triangleq P_{r}/P_{t}$ and $\quad\eta\geq 0$, where $P_{t}$ and $P_{r}$ are the transmit power and receiving power of a traveling process, respectively.
-For weak turbulent conditions, the turbulent fading coefficient is usually modeled as an LN distribution with the normalized probability density function (PDF) given by [45]
+Besides the scattering and absorption effects, the turbulence also introduces a fluctuation effect on the light irradiance [andrews2005laser].
+The turbulent fluctuation effect can be characterized by the turbulent fading coefficient $\eta$ with $\eta\triangleq P_{r}/P_{t}$ and $\eta\geq 0$, where $P_{t}$ and $P_{r}$ are the transmit power and receiving power of a traveling process, respectively.
+For weak turbulent conditions, the turbulent fading coefficient is usually modeled as an LN distribution with the normalized probability density function (PDF) given by [andrews2005laser]
 $f_{H,LN}(\eta;d)=\frac{1}{\eta\sqrt{2\pi\sigma_{I}^{2}}}\exp\left({-\frac{(\ln\eta+0.5\sigma_{I}^{2}(d))^{2}}{2\sigma_{I}^{2}(d)}}\right),$ (8)
-where $\sigma_{I}^{2}(d)=e^{\sigma_{r}^{2}(d)}-1$ and $\sigma_{r}^{2}(d)$ is the Rytov variance for a propagating distance $d$. For a plane wave, the Rytov variance is given by [45]
+where $\sigma_{I}^{2}(d)=e^{\sigma_{r}^{2}(d)}-1$ and $\sigma_{r}^{2}(d)$ is the Rytov variance for a propagating distance $d$. For a plane wave, the Rytov variance is given by [andrews2005laser]
 $\sigma_{r}^{2}(d)=1.23C_{n}^{2}k^{7/6}d^{11/6}.$ (9)
-For moderate and strong turbulent conditions, the turbulent fading coefficient is usually modeled as a GG distribution with the normalized PDF given by [45, 57]
-$f_{H,GG}(\eta;d)$ $=\frac{2[\alpha(d)\beta(d)]^{\frac{\alpha(d)+\beta(d)}{2}}\eta^{\frac{\alpha(d)+\beta(d)}{2}-1}}{\Gamma(\alpha(d))\Gamma(\beta(d))}K_{\alpha(d)-\beta(d)}\left(2\sqrt{\alpha(d)\beta(d)\eta}\right),\quad\eta\geq 0,$ (13)
-where $K_{v}(x)$ is the modified Bessel function of the second kind with parameter $v$; $\alpha(d)$ and $\beta(d)$ are related to the Rytov variance $\sigma_{r}^{2}(d)$ as [45, 57]
+For moderate and strong turbulent conditions, the turbulent fading coefficient is usually modeled as a GG distribution with the normalized PDF given by [andrews2005laser]
+$f_{H,GG}(\eta;d)$ $=\frac{2[\alpha(d)\beta(d)]^{\frac{\alpha(d)+\beta(d)}{2}}\eta^{\frac{\alpha(d)+\beta(d)}{2}-1}}{\Gamma(\alpha(d))\Gamma(\beta(d))}$ (10)
+$\quad\quad\times K_{\alpha(d)-\beta(d)}\left(2\sqrt{\alpha(d)\beta(d)\eta}\right),\quad\eta\geq 0,$
+where $K_{v}(x)$ is the modified Bessel function of the second kind with parameter $v$; $\alpha(d)$ and $\beta(d)$ are related to the Rytov variance $\sigma_{r}^{2}(d)$ as [andrews2005laser]
 $\begin{cases}\alpha(d)=\left[\exp\left(\frac{0.49\sigma_{r}^{2}(d)}{(1+1.11\sigma_{r}^{12/5})^{7/6}}\right)-1\right]^{-1},\\ \beta(d)=\left[\exp\left(\frac{0.51\sigma_{r}^{2}(d)}{(1+0.69\sigma_{r}^{12/5})^{5/6}}\right)-1\right]^{-1}.\end{cases}$ (11)
@@
 Figure 1: Geometry setting for the multiple-scattering process of NLOS UV communications
-We first introduce the geometry setting for the multiple-scattering process of NLOS UV communications, shown in Fig.
+The geometry setting for the multiple-scattering process of NLOS UV communications is shown in Fig.
 1.
 The receiver (Rx) locates at the origin $(0,0,0)$ and the transmitter (Tx) locates at the $y$-axis with coordinates $(0,r,0)$.
-The direction cosines of Tx and Rx pointing directions are denoted by $\bm{\mu}_{T}$ and $\bm{\mu}_{R}$, respectively; The zenith angle and the azimuth angle of the Tx(Rx) are denoted by $\theta_{T(R)}$ and $\phi_{T(R)}$, respectively.
+The direction cosines of Tx and Rx pointing directions are denoted by $\bm{\mu}_{T}$ and $\bm{\mu}_{R}$, respectively.
+The zenith angle and the azimuth angle of the Tx (Rx) are denoted by $\theta_{T(R)}$ and $\phi_{T(R)}$, respectively.
 The divergence angle of the Tx light beam is denoted by $\beta_{T}$ and the divergence angle of the Rx field-of-view (FOV) is denoted by $\beta_{R}$.
-We regard the initial photon emitting as the $zero$-order scattering and denote the propagating distance, scattering zenith angle and scattering azimuth angle of $i$-order scattering by $d_{i}$, $\theta_{i}$ and $\phi_{i}$, respectively.
+We regard the initial photon emitting as the zerozero-order scattering and denote the propagating distance, scattering zenith angle and scattering azimuth angle of $i$-order scattering by $d_{i}$, $\theta_{i}$ and $\phi_{i}$, respectively.
 The position of the $i$th scatter is denoted by $\bm{r}_{i}$.
@@
 We denote the set of parameters satisfying this geometrical constrain by $\mathcal{F}\triangleq\{\bm{d},\bm{\theta},\bm{\phi}|\bm{r}_{n}\text{ in FOV}\}=\{\bm{d},\bm{\theta},\bm{\phi}|\bm{r}_{n}\cdot\bm{\mu}_{R}\geq d_{n}\cos(\beta_{R}/2)\}$.
-Then the conditional receiving probability of an $n$-order scattering process for a given propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$ can be approximated as [17]
-$P_{n}(\bm{d},\bm{\theta},\bm{\phi})$ $=I_{n}(\bm{d},\bm{\theta},\bm{\phi})\left(\frac{k_{s}^{tot}}{k_{e}^{tot}}\right)^{n}\exp({-k_{e}^{tot}d_{n}})\cos\phi_{r}\min\left(1,p^{tot}(\theta_{r,n})\Omega_{r}\right),$ (15)
-where $d_{n}\triangleq\|\bm{r}_{n}\|$ and $\theta_{r,n}$ is the angle between $\bm{\mu}_{n-1}$ and $-\bm{r}_{n}$; $\phi_{r}$ is the angle between $\bm{\mu}_{R}$ and $\bm{r}_{n}$; $\Omega_{r}$ is the solid angle formed by $\bm{r}_{n}$ and the receiving area $A_{r}$; and we have $\Omega_{r}\approx 2\pi(1-d_{n}/\sqrt{d_{n}^{2}+r_{A}^{2}})$, where $r_{A}=\sqrt{A_{r}/\pi}$ is the radius of the receiving area; $I_{n}(\bm{d},\bm{\theta},\bm{\phi})$ is an indicator function; and we have $I_{n}(\bm{d},\bm{\theta},\bm{\phi})=1$ and $I_{n}(\bm{d},\bm{\theta},\bm{\phi})=0$ when the $n$th scatter locates within and without the FOV, respectively.
+Then the conditional receiving probability of an $n$-order scattering process for a given propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$ can be approximated as [yuan2019monte]
+$P_{n}(\bm{d},\bm{\theta},\bm{\phi})$ $=I_{n}(\bm{d},\bm{\theta},\bm{\phi})\left(\frac{k_{s}^{tot}}{k_{e}^{tot}}\right)^{n}\exp({-k_{e}^{tot}d_{n}})$ (12)
+$\quad\quad\times\cos\phi_{r}\min\left(1,p^{tot}(\theta_{r,n})\Omega_{r}\right),$
+where $d_{n}\triangleq\|\bm{r}_{n}\|$ and $\theta_{r,n}$ is the angle between $\bm{\mu}_{n-1}$ and $\bm{r}_{n}$; $\phi_{r}$ is the angle between $\bm{\mu}_{R}$ and $\bm{r}_{n}$, as shown in Fig.
+1; $\Omega_{r}$ is the solid angle formed by $\bm{r}_{n}$ and the receiving area $A_{r}$; and we have $\Omega_{r}\approx 2\pi(1-d_{n}/\sqrt{d_{n}^{2}+r_{A}^{2}})$, where $r_{A}=\sqrt{A_{r}/\pi}$ is the radius of the receiving area; and $I_{n}(\bm{d},\bm{\theta},\bm{\phi})$ is an indicator function.
+The indicator function $I_{n}(\bm{d},\bm{\theta},\bm{\phi})$ is defined as
+$I_{n}(\bm{d},\bm{\theta},\bm{\phi})=\begin{cases}1,&\text{$\bm{r_{n}}$ locates in FOV},\\ 0,&\text{otherwise}.\end{cases}$ (13)
 Then the conditional receiving probability can be rewritten as
 $P_{n}(\bm{d},\bm{\theta},\bm{\phi})=\begin{cases}P_{d,n}(\bm{d},\bm{\theta},\bm{\phi}),\quad\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F},\\ 0,\quad\{\bm{d},\bm{\theta},\bm{\phi}\}\notin\mathcal{F},\end{cases}$ (14)
-where $P_{d,n}(\bm{d},\bm{\theta},\bm{\phi})\triangleq\left(\frac{k_{s}^{tot}}{k_{e}^{tot}}\right)^{n}\exp({-k_{e}^{tot}d_{n}})\cos\phi_{r}\min\left(1,\right.$ $\left.p^{tot}(\theta_{r,n})\Omega_{r}\right)$.
+where $P_{d,n}(\bm{d},\bm{\theta},\bm{\phi})\triangleq\left(\frac{k_{s}^{tot}}{k_{e}^{tot}}\right)^{n}\exp({-k_{e}^{tot}d_{n}})\cos\phi_{r}\min\left(1,\right$.
+$\left.p^{tot}(\theta_{r,n})\Omega_{r}\right)$.
 Then the receiving probability can be obtained by averaging $\{\bm{d},\bm{\theta},\bm{\phi}\}$ out, i.e.,
-$P_{n}$ $=\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}P_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi},$ (17)
-where $f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})=\prod_{i=0}^{n-1}f_{D}(d_{i})f_{\Theta}(\theta_{i})f_{\Phi}(\phi_{i})$ is the joint PDF of $\{\bm{d},\bm{\theta},\bm{\phi}\}$.
+$P_{n}$ $=\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}P_{n}(\bm{d},\bm{\theta},\bm{\phi})f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi},$ (15)
+where $f(\bm{d},\bm{\theta},\bm{\phi})=\prod_{i=0}^{n-1}f_{D}(d_{i})f_{\Theta}(\theta_{i})f_{\Phi}(\phi_{i})$ is the joint PDF of $\{\bm{d},\bm{\theta},\bm{\phi}\}$.
 We have used the fact that $\bm{d}$, $\bm{\theta}$, and $\bm{\phi}$ are independent from each other.
-For the propagating distance $d_{i}$ with $i\in\{0,1,\cdots,n-1\}$, we have $f_{D}(d_{i})=k_{e}^{tot}e^{-k_{e}^{tot}d_{i}}$ with $0\leq d_{i}\leq\infty$ [17].
-For the scattering angle $\theta_{i}$ with $i=0$, we have $f_{\Theta}(\theta_{0})=\frac{\sin\theta_{0}}{1-\cos({\beta_{T}}/{2})}$ with $0\leq\theta_{0}\leq{\beta_{T}}/{2}$ when a uniform light source is assumed [17]; and for the scattering angle $\theta_{i}$ with $i\in\{1,2,\cdots,n-1\}$, we have $f_{\Theta}(\theta_{i})=2\pi p^{tot}(\theta_{i})\sin\theta_{i}$ with $0\leq\theta_{i}\leq\pi$ [17].
-For the scattering azimuth angle $\phi_{i}$ with $i\in\{0,1,\cdots,n-1\}$, we have $f_{\Phi}(\phi_{i})=\frac{1}{2\pi}$ with $0\leq\phi_{i}\leq 2\pi$ [17].
+For the propagating distance $d_{i}$ with $i\in\{0,1,\cdots,n-1\}$, we have $f_{D}(d_{i})=k_{e}^{tot}e^{-k_{e}^{tot}d_{i}}$ with $0\leq d_{i}\leq\infty$ [yuan2019monte].
+For the scattering angle $\theta_{i}$ with $i=0$, we have $f_{\Theta}(\theta_{0})=\frac{\sin\theta_{0}}{1-\cos({\beta_{T}}/{2})}$ with $0\leq\theta_{0}\leq{\beta_{T}}/{2}$ when a uniform light source is assumed [yuan2019monte]; and for the scattering angle $\theta_{i}$ with $i\in\{1,2,\cdots,n-1\}$, we have $f_{\Theta}(\theta_{i})=2\pi p^{tot}(\theta_{i})\sin\theta_{i}$ with $0\leq\theta_{i}\leq\pi$ [yuan2019monte].
+For the scattering azimuth angle $\phi_{i}$ with $i\in\{0,1,\cdots,n-1\}$, we have $f_{\Phi}(\phi_{i})=\frac{1}{2\pi}$ with $0\leq\phi_{i}\leq 2\pi$ [yuan2019monte].
 Without loss of generality, we assume that a unit light power is transmitted from the light source. Then the average receiving power equals the receiving probability $P_{n}$.
@@
 The conditional receiving probability without turbulent fluctuation for a given propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$ is $P_{n}(\bm{d},\bm{\theta},\bm{\phi})$ given in (12).
-Then the conditional receiving probability under turbulent fluctuation effect for a given set of turbulent fading coefficients $\bm{\eta}\triangleq[\eta_{0},\eta_{1},\cdots,\eta_{n}]^{\text{T}}$ and a propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$ can be expressed as
-$P_{tur,n}(\bm{\eta},\bm{d},\bm{\theta},\bm{\phi})=\prod_{i=0}^{n}\eta_{i}P_{n}(\bm{d},\bm{\theta},\bm{\phi}).$ (18)
+Then the conditional receiving probability under turbulent fluctuation effect for a given set of turbulent fading coefficients $\bm{\eta}\triangleq[\eta_{0},\eta_{1},\cdots,\eta_{n}]^{\text{T}}$ under a propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$ can be expressed as
+$P_{tur,n}(\bm{\eta},\bm{d},\bm{\theta},\bm{\phi})=P_{n}(\bm{d},\bm{\theta},\bm{\phi})\prod_{i=0}^{n}\eta_{i}.$ (16)
 Then the receiving probability under turbulent fluctuation for a given propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$ can be obtained by averaging $\bm{\eta}$ out, i.e.,
-$P_{tur,n}(\bm{d},\bm{\theta},\bm{\phi})$ $=\int_{\bm{\eta}}P_{tur,n}(\bm{\eta},\bm{d},\bm{\theta},\bm{\phi})f_{H|D,\Theta,\Phi}(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{\eta},$ (19)
-where $f_{H|D,\Theta,\Phi}(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})$ is the joint PDF of $\{\eta_{0},\eta_{1},\cdots,\eta_{n}\}$ conditioned on a given propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$.
+$P_{tur,n}(\bm{d},\bm{\theta},\bm{\phi})$ $=\int_{\bm{\eta}}P_{tur,n}(\bm{\eta},\bm{d},\bm{\theta},\bm{\phi})f(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{\eta},$ (17)
+where $f(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})$ is the joint PDF of $\{\eta_{0},\eta_{1},\cdots,\eta_{n}\}$ conditioned on a given propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$.
 For a given propagating path, $\eta_{0},\eta_{1},\cdots,\eta_{n}$ are assumed as independent variables.
 Then we have
-$f_{H|D,\Theta,\Phi}(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})=\prod_{i=0}^{n}f_{H}(\eta_{i};d_{i}),$ (20)
-where $f_{H}(\eta;d)=f_{H,LN}(\eta;d)$ is given in (11) for LN distribution and $f_{H}(\eta;d)=f_{H,GG}(\eta;d)$ is given in (13) for GG distribution.
+$f(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})=\prod_{i=0}^{n}f(\eta_{i};\sigma_{I}^{2}(d_{i})),$ (18)
+where $f(\eta;d)=f_{H,LN}(\eta;d)$ is given in (8) for LN distribution and $f(\eta;d)=f_{H,GG}(\eta;d)$ is given in (10) for GG distribution.
 By substituting (16) and (18) into (17), we can obtain
-$P_{tur,n}(\bm{d},\bm{\theta},\bm{\phi})$ $=P_{n}(\bm{d},\bm{\theta},\bm{\phi})\prod_{i=0}^{n}\left[\int_{\eta_{i}}\eta_{i}f_{H}(\eta_{i};d_{i})\mathrm{d}\eta_{i}\right]=P_{n}(\bm{d},\bm{\theta},\bm{\phi}),$ (21)
-where we have used the property that $\int_{\eta_{i}}\eta_{i}f_{H}(\eta_{i};d_{i})\mathrm{d}\eta_{i}=1$ for $i=0,1,\cdots,n$.
+$P_{tur,n}(\bm{d},\bm{\theta},\bm{\phi})$ $=P_{n}(\bm{d},\bm{\theta},\bm{\phi})\prod_{i=0}^{n}\left[\int_{\eta_{i}}\eta_{i}f(\eta_{i};\sigma_{I}^{2}(d_{i}))\mathrm{d}\eta_{i}\right]$ (19)
+$=P_{n}(\bm{d},\bm{\theta},\bm{\phi}),$
+where we have used the property $\int_{\eta_{i}}\eta_{i}f(\eta_{i};\sigma_{I}^{2}(d_{i}))\mathrm{d}\eta_{i}=1$ for $i=0,1,\cdots,n$.
 Then the receiving probability considering the turbulent fluctuation effect can be obtained by averaging $\{\bm{d},\bm{\theta},\bm{\phi}\}$ out, i.e.,
-$P_{tur,n}$ $=\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}P_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}=P_{n}.$ (22)
+$P_{tur,n}$ $=\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}P_{n}(\bm{d},\bm{\theta},\bm{\phi})f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}=P_{n}.$ (20)
 From (20), we can see that $P_{tur,n}=P_{n}$, which indicates that the average receiving power considering the turbulent fluctuation equals the average receiving power ignoring the turbulent fluctuation.
 Therefore, in practical implementations, we can estimate the average receiving power by solving the integration in (15).
-IV Estimating Turbulent Channels Using MCI Approach
+IV Estimating Turbulent Channels by Using an MCI Approach
 IV-A Estimating the Average Receiving Power
-The integration in (17) can be estimated by using the Monte-Carlo methods.
-In this paper, we use the MCI approach to estimate the average receiving power in (17).
-Compared with another widely used Monte-Carlo simulation (MCS) approach [13, 14], the MCI approach [15, 16, 17, 18, 19] enjoys a more flexible sampling function, resulting in a higher computational efficiency.
-Specifically, we can rewrite (17) as
+The integration in (15) can be estimated by using an MCI approach.
+Compared with another widely used MCS approach [ding2009modeling, drost2011uv], the MCI approach [yuan2019importance, ding2010path, yuan2019monte, shan2020modeling, shen2020improved] enjoys a more flexible sampling function, resulting in a higher computational efficiency.
+Besides, as we will demonstrate later in Section IV-B, the MCI approach can provide a more interpretable logic for estimating the turbulent fluctuation.
+Following the MCI procedure given in [yuan2019monte], we can rewrite (15) as
 $P_{n}=\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}g_{n}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi},$ (21)
-where $g_{n}(\bm{d},\bm{\theta},\bm{\phi})$ is defined as $g_{n}(\bm{d},\bm{\theta},\bm{\phi})\triangleq P_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})$, and where $P_{n}(\bm{d},\bm{\theta},\bm{\phi})$ is given in (16).
+where $g_{n}(\bm{d},\bm{\theta},\bm{\phi})$ is defined as $g_{n}(\bm{d},\bm{\theta},\bm{\phi})\triangleq P_{n}(\bm{d},\bm{\theta},\bm{\phi})f(\bm{d},\bm{\theta},\bm{\phi})$, and where $P_{n}(\bm{d},\bm{\theta},\bm{\phi})$ is given in (14).
 Now we can choose a sampling PDF $f_{n}(\bm{d},\bm{\theta},\bm{\phi})$ and further rewrite the receiving power $P_{n}$ in (21) as
-$P_{n}$ $=\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}O_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{n}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}=\text{E}[O_{n}(\bm{d},\bm{\theta},\bm{\phi})],$ (24)
-where $O_{n}(\bm{d},\bm{\theta},\bm{\phi})\triangleq g_{n}(\bm{d},\bm{\theta},\bm{\phi})/f_{tur,n}(\bm{d},\bm{\theta},\bm{\phi})$ is called the objective function of MCI approach; and $\text{E}[x]$ denotes the expectation of $x$.
-Therefore, we have expressed the receiving power $P_{n}$ as the expectation of the objective function $O_{n}(\bm{d},\bm{\theta},\bm{\phi})$ when the joint PDF of $\{\bm{d},\bm{\theta},\bm{\phi}\}$ is the sampling PDF $f_{n}(\bm{d},\bm{\theta},\bm{\phi})$.
-Different sampling PDF $f_{n}(\bm{d},\bm{\theta},\bm{\phi})$ can result in different computational efficiency of MCI approach [15, 16, 17].
-Here we adopt the so-called important sampling method [15], which enjoys the same convergent speed as the MCS approach.
+$P_{n}$ $=\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}O_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{n}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}$ (22)
+$=\text{E}[O_{n}(\bm{d},\bm{\theta},\bm{\phi})],$
+where $O_{n}(\bm{d},\bm{\theta},\bm{\phi})\triangleq g_{n}(\bm{d},\bm{\theta},\bm{\phi})/f_{n}(\bm{d},\bm{\theta},\bm{\phi})$ is called the objective function of MCI approach; and $\text{E}[x]$ denotes the expectation of $x$.
+Therefore, we have expressed the receiving power $P_{n}$ as the expectation of the objective function $O_{n}(\bm{d},\bm{\theta},\bm{\phi})$ when a sampling PDF $f_{n}(\bm{d},\bm{\theta},\bm{\phi})$ is used to sample $\{\bm{d},\bm{\theta},\bm{\phi}\}$.
+Different sampling PDF $f_{n}(\bm{d},\bm{\theta},\bm{\phi})$ can result in different computational efficiency of MCI approach [yuan2019importance, ding2010path, yuan2019monte].
+Here we adopt the so-called important sampling method [yuan2019importance], which enjoys the same convergent speed as the MCS approach but a faster calculating speed.
 The sampling PDF of the important sampling is given by
-$f_{n}(\bm{d},\bm{\theta},\bm{\phi})=f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})=\prod_{i=0}^{n-1}f_{D}(d_{i})f_{\Theta}(\theta_{i})f_{\Phi}(\phi_{i}).$ (25)
+$f_{n}(\bm{d},\bm{\theta},\bm{\phi})=f(\bm{d},\bm{\theta},\bm{\phi})=\prod_{i=0}^{n-1}f_{D}(d_{i})f_{\Theta}(\theta_{i})f_{\Phi}(\phi_{i}).$ (23)
 According to the sampling PDF $f_{n}(\bm{d},\bm{\theta},\bm{\phi})$ in (23), we can obtain the sampling functions for generating the propagating distance, the scattering angle, and the scattering azimuth angle as
@@
 $\phi_{i}=2\pi\times\text{rand}(1),\quad i=0,1,\cdots,n-1,$ (26)
-where $\text{rand}(1)$ is a random number between 0 and 1; $F_{\Theta}^{-1}(\theta_{s})$ is the inverse function of the cumulative distribution function of the scattering angle, i.e., $F_{\Theta}(\theta_{s})=\int_{0}^{\theta_{s}}f_{\Theta}(\theta)\mathrm{d}\theta$, which can be obtained as
-$F_{\Theta}(\theta_{s})$ $=C_{1}\cos\theta_{s}+C_{2}\cos^{3}\theta_{s}+C_{3}+C_{4}(C_{5}-C_{6}\cos\theta_{s})^{-1/2}+\frac{C_{7}(1-\cos\theta_{s})}{2+C_{8}(1-\cos\theta_{s})},$ (29)
-where $C_{1}=-\frac{3(1+3\gamma)}{8(1+2\gamma)}\frac{k_{s}^{ray}}{k_{s}^{tot}}+\frac{f(1-g^{2})}{4(1+g^{2})^{3/2}}\frac{k_{s}^{mie}}{k_{s}^{tot}}$, $C_{2}=-\frac{1-\gamma}{8(1+2\gamma)}\frac{k_{s}^{ray}}{k_{s}^{tot}}-\frac{f(1-g^{2})}{4(1+g^{2})^{3/2}}\frac{k_{s}^{mie}}{k_{s}^{tot}}$, $C_{3}=\frac{1}{2}\frac{k_{s}^{ray}}{k_{s}^{tot}}+\frac{1+g}{2g}\frac{k_{s}^{mie}}{k_{s}^{tot}}$, $C_{4}=-\frac{1-g^{2}}{2g}\frac{k_{s}^{mie}}{k_{s}^{tot}}$, $C_{5}=1+g^{2}$, $C_{6}=2g$, $C_{7}=(1+4k^{2}d_{0}^{2})\frac{k_{s}^{tur}}{k_{s}^{tot}}$, and $C_{8}=4k^{2}d_{0}^{2}$.
+where $\text{rand}(1)$ is a random number between 0 and 1; $F_{\Theta}^{-1}(\theta_{s})$ denotes the inverse cumulative distribution function of the scattering angle $\theta_{s}$.
 For $i=1,2,\cdots,n-1$, the scattering angle $\theta_{i}$ can be solved by using numerical methods, e.g., Newton’s bisection search.
 Then the corresponding objective function becomes
-$O_{n}(\bm{d},\bm{\theta},\bm{\phi})$ $=P_{n}(\bm{d},\bm{\theta},\bm{\phi})=\begin{cases}P_{d,n},\quad\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F},\\ 0,\quad\{\bm{d},\bm{\theta},\bm{\phi}\}\notin\mathcal{F}.\end{cases}$ (30)
+[…]
 According to the law of large numbers, we can estimate the expectation $P_{n}$ using the average of sampling values for the objective function.
@@
 $\bm{s}_{n}^{m}\triangleq[d_{0}^{m},\theta_{0}^{m},\phi_{0}^{m},\cdots,d_{n-1}^{m},\theta_{n-1}^{m},\phi_{n-1}^{m}]^{\text{T}}.$ (28)
-Then we can approximate the receiving power when $M$ is large as
+Then when $M$ is large, we can approximate the receiving power as
 $P_{n}=\frac{1}{M}\sum_{m=1}^{M}O_{n}(\bm{s}_{n}^{m}).$ (29)
@@
 $P_{tot}=\sum_{n=1}^{N}P_{n}=\frac{1}{M}\sum_{m=1}^{M}\sum_{n=1}^{N}O_{n}(\bm{s}_{n}^{m}).$ (30)
-IV-B Estimating the Channel Impulse Response
-For each sampling point $\bm{s}_{n}^{m}$, we can calculating the corresponding receiving time as
-$t_{n}^{m}=\frac{\sum_{i=0}^{n}d_{i}^{m}}{c},$ (34)
-where $c$ is the light speed in the atmosphere. Then according to the result in [17], we can obtain the CIR of the multiple-scattering of UV communications in the following steps:
-  • Step 1: Classify the objective function value $O_{n}(\bm{s}_{n}^{m})$ into $L$ different time slots $\{(0,\Delta t)$, $(\Delta t,2\Delta t),\cdots,((L-1)\Delta t,L\Delta t)\}$ according to its receiving time $t_{n}^{m}$, where $\Delta t$ is a pre-chosen time interval;
-  • Step 2: Add up objective function values in each time slot and divide it by the number of sampling points $M$.
-Then we obtain the receiving probability sequences $[{P}_{n}(\Delta t),{P}_{n}(2\Delta t)$, $\cdots,{P}_{n}(L\Delta t)]^{\text{T}}$;
-  • Step 3: The CIR is obtained by normalizing the sequences $[{P}_{n}(\Delta t),{P}_{n}(2\Delta t),\!\cdots,\!{P}_{n}(L\Delta t)]^{\text{T}}$ using the receiving time interval $\Delta t$ and the receiving area $A_{r}$.
-Therefore, the value of CIR in a unit of irradiance at $t=l\Delta t$ is given by
-    $h_{n}(l\Delta t)=\frac{P_{n}(l\Delta t)}{\Delta_{t}A_{r}},\quad l=1,2\cdots,L.$ (35)
-Then the total CIR at $t=l\Delta t$ over $N$ scattering orders can be obtained as
-$h_{tot}(l\Delta t)=\sum_{n=1}^{N}\frac{P_{n}(l\Delta t)}{\Delta_{T}A_{r}},\quad l=1,2\cdots,L.$ (36)
-IV-C Estimating the Turbulent Fluctuation
+IV-B Estimating the Turbulent Fluctuation
 It is challenging to estimate the turbulent fluctuation of the receiving power by using Monte-Carlo based methods.
-This is because the Monte-Carlo process itself will introduce a fluctuation effect on the calculated average receiving power [17].
-In existing literature [52], the turbulent fluctuation was estimated by simply calculating the variance of the average receiving power $P_{tur,n}$, i.e, $\text{Var}[P_{tur,n}]$, obtained from the Monte-Carlo models.
+This is because the Monte-Carlo process itself will introduce a fluctuation effect on the calculated average receiving power [yuan2019monte].
+In existing literature [wang2013characteristics], the turbulent fluctuation was estimated by simply calculating the variance of the average receiving power $P_{tur,n}$, i.e, $\text{Var}[P_{tur,n}]$, obtained from the Monte-Carlo models.
 However, because $P_{tur,n}=P_{n}$, using the expression in (29), we can obtain
@@
 where we have used the fact that all sampling points $\{\bm{s}_{n}^{1},\bm{s}_{n}^{2},\cdots,\bm{s}_{n}^{M}\}$ are independent from each other and $\text{Var}[O_{n}(\bm{s}_{n}^{1})]=\text{Var}[O_{n}(\bm{s}_{n}^{2})]\cdots=\text{Var}[O_{n}(\bm{s}_{n}^{M})]=\text{Var}[O_{n}(\bm{d},\bm{\theta},\bm{\phi})]$ is the variance of the objective function.
-We can see that the variance $\text{Var}[P_{tur,n}]$ depends on both $\text{Var}[O_{n}(\bm{d},\bm{\theta},\bm{\phi})]$ and $M$.
+From eq.
+(31), we can see that the variance $\text{Var}[P_{tur,n}]$ depends on both $\text{Var}[O_{n}(\bm{d},\bm{\theta},\bm{\phi})]$ and $M$.
 Besides, we have $\text{Var}[P_{tur,n}]\to 0$ when the number of sampling points $M\to\infty$, which means the calculated variance can be arbitrarily small as long as $M$ is large enough.
-Whereas, larger sampling points $M$ in the Monte-Carlo process only reduce the fluctuation caused by randomness, leading to more stable and precise results.
-Therefore, any attempt to calculate the turbulent fluctuation by estimating $\text{Var}[P_{tur,n}]$ will fail due to the fact that $\text{Var}[P_{tur,n}]$ depends on $M$.
-Here we present an improved estimating method for the turbulent fluctuation, which is independent of $M$.
-By substituting (16) into (18), we can obtain
-$P_{tur,n}(\bm{\eta},\bm{d},\bm{\theta},\bm{\phi})=\begin{cases}\prod_{i=0}^{n}\eta_{i}P_{d,n}(\bm{d},\bm{\theta},\bm{\phi}),\quad\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F},\\ 0,\quad\{\bm{d},\bm{\theta},\bm{\phi}\}\notin\mathcal{F},\end{cases}$ (38)
+This is because a larger number of sampling points $M$ in the Monte-Carlo process will reduce the fluctuation caused by randomness, leading to more stable and precise results.
+Therefore, the attempt to calculate the turbulent fluctuation by estimating $\text{Var}[P_{tur,n}]$ becomes unreliable due to the fact that $\text{Var}[P_{tur,n}]$ depends on $M$.
+Here we present an improved method based on an MCI approach for estimating the turbulent fluctuation, which is more interpretable compared with the existing MCS approach and is independent of $M$ when $M$ is large.
+We first substitute (14) into (16) and obtain the instantaneous receiving probability as
+$P_{tur,n}(\bm{\eta},\bm{d},\bm{\theta},\bm{\phi})$ (32)
+[…]
 which can be regarded as a function of random variables $\bm{\eta}$, $\bm{d}$, $\bm{\theta}$, and $\bm{\phi}$.
@@
 To characterize the turbulent fluctuation effect only, we can average $\{\bm{d},\bm{\theta},\bm{\phi}\}$ out and obtain the instantaneous receiving power conditioned on turbulent fading coefficients as
-$P_{tur,n}(\bm{\eta})$ $=\!\prod_{i=0}^{n}\eta_{i}\!\!\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}\!P_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}=\!{\prod_{i=0}^{n}\eta_{i}}P_{n}.$ (39)
-Then we can define an equivalent turbulent fading coefficient $\eta_{eq,n}\triangleq P_{tur,n}(\bm{\eta})/P_{n}$ on the average receiving power $P_{n}$ such that the instantaneous receiving power $P_{tur,n}(\bm{\eta})=\eta_{eq,n}P_{n}$.
-Obviously, we have
+$P_{tur,n}(\bm{\eta})$ $=\!\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}P_{n}(\bm{d},\bm{\theta},\bm{\phi})\!\prod_{i=0}^{n}\eta_{i}\!f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}$ (33)
+$=\!{\prod_{i=0}^{n}\eta_{i}}P_{n}.$
+Then we can define an equivalent turbulent fading coefficient as
+$\eta_{eq,n}\triangleq\frac{P_{tur,n}(\bm{\eta})}{P_{n}}$ (34)
+on the average receiving power $P_{n}$ such that the instantaneous receiving power $P_{tur,n}(\bm{\eta})=\eta_{eq,n}P_{n}$. Obviously, we have
 $\eta_{eq,n}=\prod_{i=0}^{n}\eta_{i}.$ (35)
 One can easily verify that $\text{E}[\eta_{eq,n}]=1$. Then the turbulent fluctuation for the $n$th scattering order can be fully characterized by the turbulent variance
-$\sigma_{tur,n}^{2}$ $\triangleq\text{Var}[\eta_{eq,n}]=\int_{\bm{\eta}}(\eta_{eq,n}-1)^{2}f_{H}(\bm{\eta})\mathrm{d}\bm{\eta},$ (41)
-where $f_{H}(\bm{\eta})$ is the joint PDF of $\{\eta_{0},\eta_{1},\cdots,\eta_{n}\}$.
+$\sigma_{tur,n}^{2}$ $\triangleq\text{Var}[\eta_{eq,n}]=\int_{\bm{\eta}}(\eta_{eq,n}-1)^{2}f(\bm{\eta})\mathrm{d}\bm{\eta},$ (36)
+where $f(\bm{\eta})$ is the joint PDF of $\{\eta_{0},\eta_{1},\cdots,\eta_{n}\}$.
 From (32) we can see that those propagating paths with $\{\bm{d},\bm{\theta},\bm{\phi}\}\notin\mathcal{F}$ contribute zero to the turbulent fluctuation effect.
 Therefore, we should restrict the variance on the set $\mathcal{F}$.
-Then the joint PDF $f_{H}(\bm{\eta})$ can be obtained by averaging $\{\bm{d},\bm{\theta},\bm{\phi}\}$ out of the conditional PDF $f_{H|D,\Theta,\Phi}(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})$ in (20) on the set $\mathcal{F}$, i.e.,
-$f_{H}(\bm{\eta})$ $\!=\!\int_{\mathcal{F}}f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi}|\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F})f_{H|D,\Theta,\Phi}(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi},$ (42)
-where $f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi}|\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F})$ is the joint PDF of $\{\bm{d},\bm{\theta},\bm{\phi}\}$ on the set $\mathcal{F}$; and we have
-$f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi}|\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F})$ $=\frac{f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})}{\text{Pr}[\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F}]},$ (43)
+Then the joint PDF $f(\bm{\eta})$ can be obtained by averaging $\{\bm{d},\bm{\theta},\bm{\phi}\}$ out of the conditional PDF $f(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})$ in (18) on the set $\mathcal{F}$, i.e.,
+$f_{H}(\bm{\eta})$ $\!=\!\int_{\mathcal{F}}f(\bm{d},\bm{\theta},\bm{\phi}|\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F})$ (37)
+$\quad\quad\times f(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi},$
+where $f(\bm{d},\bm{\theta},\bm{\phi}|\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F})$ is the joint PDF of $\{\bm{d},\bm{\theta},\bm{\phi}\}$ on the set $\mathcal{F}$; and we have
+$f(\bm{d},\bm{\theta},\bm{\phi}|\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F})$ $=\frac{f(\bm{d},\bm{\theta},\bm{\phi})}{\text{Pr}[\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F}]},$ (38)
 where $\text{Pr}[\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F}]$ is the probability of $\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F}$ given by
-$\text{Pr}[\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F}]=\int_{\mathcal{F}}f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}.$ (44)
+$\text{Pr}[\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F}]=\int_{\mathcal{F}}f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}.$ (39)
 By substituting (18), (38), and (39) into (37), we can obtain
-$f_{H}(\bm{\eta})$ $=\frac{\int_{\mathcal{F}}\prod_{i=0}^{n}f_{H}(\eta_{i};d_{i})f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\mathcal{F}}f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}$ (45)
-$=\frac{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}\prod_{i=0}^{n}f_{H}(\eta_{i};d_{i})I_{n}f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}I_{n}f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}.$
+$f_{H}(\bm{\eta})$ $=\frac{\int_{\mathcal{F}}\prod_{i=0}^{n}f(\eta_{i};d_{i})f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\mathcal{F}}f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}$ (40)
+$=\frac{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}\prod_{i=0}^{n}f(\eta_{i};d_{i})I_{n}f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}I_{n}f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}.$
 Then by substituting (35) and (40) into (36), we can obtain the turbulent variance as
-$\sigma_{tur,n}^{2}$ $\!=\!\frac{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}(\prod_{i=0}^{n}M_{2}(d_{i})-1)I_{n}f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}I_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}},$ (46)
+$\sigma_{tur,n}^{2}$ $\!=\!\frac{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}(\prod_{i=0}^{n}M_{2}(d_{i})-1)I_{n}f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}I_{n}f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}},$ (41)
 where $M_{2}(d_{i})$ is the second-order moment of the turbulent fading coefficient $\eta_{i}(d_{i})$.
@@
 Then using the same sampling points $\{\bm{s}_{n}^{1},\bm{s}_{n}^{2},\cdots,\bm{s}_{n}^{M}\}$, we can estimate the turbulent variance $\sigma_{tur,n}^{2}$ as
-$\sigma_{tur,n}^{2}$ $=\frac{\frac{1}{M}\sum_{m=1}^{M}\left[\prod_{i=0}^{n}M_{2}(d_{i}^{m})-1\right]I_{n}(\bm{s}_{n}^{m})}{\frac{1}{M}\sum_{m=0}^{M}I_{n}(\bm{s}_{n}^{m})}=\frac{1}{Count}\sum_{m=1}^{M}\left[\prod_{i=0}^{n}M_{2}(d_{i}^{m})-1\right]I_{n}(\bm{s}_{n}^{m}),$ (47)
+$\sigma_{tur,n}^{2}$ $=\frac{\frac{1}{M}\sum_{m=1}^{M}\left[\prod_{i=0}^{n}M_{2}(d_{i}^{m})-1\right]I_{n}(\bm{s}_{n}^{m})}{\frac{1}{M}\sum_{m=0}^{M}I_{n}(\bm{s}_{n}^{m})}$ (42)
+$=\frac{1}{Count}\sum_{m=1}^{M}\left[\prod_{i=0}^{n}M_{2}(d_{i}^{m})-1\right]I_{n}(\bm{s}_{n}^{m}),$
 where $Count\triangleq\sum_{m=0}^{M}I_{n}(\bm{s}_{n}^{m})$ is the number of sampling points satisfying $\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F}$.
-Similar to the $n$th scattering order, the total turbulent fluctuation effect on the total receiving power of $N$ scattering orders can be characterized by defining an equivalent turbulent fading coefficient $\eta_{eq}\triangleq\sum_{n=1}^{N}P_{tur,n}(\bm{\eta})/\sum_{n=1}^{N}P_{n}$.
+Similar to the $n$th scattering order, the total turbulent fluctuation effect on the total receiving power of $N$ scattering orders can be characterized by defining an equivalent turbulent fading coefficient as
+$\eta_{eq}\triangleq\frac{\sum_{n=1}^{N}P_{tur,n}(\bm{\eta})}{\sum_{n=1}^{N}P_{n}},$ (43)
+which allows us to treat the complex multiple-scattering channel as a single equivalent link with a unified fading parameter and simplify the characterization of turbulent fluctuation effects.
 Using the relations $P_{tur,n}(\bm{\eta})=\eta_{eq,n}P_{n}$ and $P_{tot}=\sum_{n=1}^{N}P_{n}$, we can obtain
@@
 where $P_{n}$, $P_{tot}$, and $\sigma_{tur,n}^{2}$ are given by (29), (30), and (42), respectively.
-IV-D Estimating the Distribution of Turbulent Fading Coefficient
-It is challenging to derive the distribution of the turbulent fading coefficient for a general turbulent fading model.
-However, when the turbulence is weak and an LN fading is considered, then we can further derive the distributions of the equivalent turbulent fading coefficients $\eta_{eq,n}$ and $\eta_{eq}$.
+IV-C Estimating the Distribution of Turbulent Fading Coefficient
+Now we further derive the distributions of the equivalent turbulent fading coefficients $\eta_{eq,n}$ and $\eta_{eq}$.
 For a given propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$, the equivalent turbulent fading coefficient $\eta_{eq,n}$ can be obtained from (35) as
@@
 $\ln\eta_{eq,n}(\bm{d},\bm{\theta},\bm{\phi})=\sum_{i=0}^{n}\ln\eta_{i}(\bm{d},\bm{\theta},\bm{\phi}).$ (47)
-Because $\eta_{i}(\bm{d},\bm{\theta},\bm{\phi})$ satisfies an LN distribution with parameter $\sigma_{I}^{2}(d_{i})$ and $\{\eta_{0},\eta_{1},\cdots,\eta_{n}$ are independent variables, from (51) we can see that $\eta_{eq,n}(\bm{d},\bm{\theta},\bm{\phi})$ also satisfies an LN distribution with parameter $\sigma_{eq,n}^{2}(\bm{d},\bm{\theta},\bm{\phi})=\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i})$.
+When $\eta_{i}(\bm{d},\bm{\theta},\bm{\phi})$ satisfies an LN distribution with parameter $\sigma_{I}^{2}(d_{i})$ and $\eta_{0},\eta_{1},\cdots,\eta_{n}$ are independent variables, from (47) we can see that $\eta_{eq,n}(\bm{d},\bm{\theta},\bm{\phi})$ also satisfies an LN distribution with parameter $\sigma_{eq,n}^{2}(\bm{d},\bm{\theta},\bm{\phi})=\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i})$.
 Therefore, the conditional PDF of $\eta_{eq,n}(\bm{d},\bm{\theta},\bm{\phi})$ is given by
-$f_{H,n}(\eta_{eq,n}|\bm{d},\bm{\theta},\bm{\phi})$ $=\frac{1}{\eta_{eq,n}\sqrt{2\pi\sigma_{eq,n}^{2}(\bm{d},\bm{\theta},\bm{\phi})}}\exp\left(\!{-\frac{(\ln\eta_{eq,n}+0.5\sigma_{eq,n}^{2}(\bm{d},\bm{\theta},\bm{\phi}))^{2}}{2\sigma_{eq,n}^{2}(\bm{d},\bm{\theta},\bm{\phi})}}\!\right).$ (52)
-Then the unconditional PDF for the equivalent turbulent fading coefficient $\eta_{eq,n}$ can be obtained by averaging $\{\bm{d},\bm{\theta},\bm{\phi}\}$ on $\mathcal{F}$ out, which is given by
-$f_{H,n}(\eta_{eq,n})$ $=\frac{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}\frac{I_{n}(\bm{d},\bm{\theta},\bm{\phi})}{\eta_{eq,n}\sqrt{2\pi\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i})}}\exp\left({-\frac{(\ln\eta_{eq,n}+0.5\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i}))^{2}}{2\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i})}}\right)f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}I_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}.$ (53)
-Then we can also use the MCI approach to estimate $f_{H}(\eta_{eq,n})$ using the same sampling PDF given in (25).
-Using the same sampling points $\{\bm{s}_{n}^{1},\bm{s}_{n}^{2},\cdots,\bm{s}_{n}^{M}\}$, we can estimate the PDF of equivalent turbulent fading coefficient $f_{H}(\eta_{eq,n})$ as
-$f_{H,n}(\eta_{eq,n})$ $=\frac{1}{Count}\sum_{m=1}^{M}\frac{I_{n}(\bm{s}_{n}^{m})}{\eta_{eq,n}\sqrt{2\pi\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i}^{m})}}\exp\left({-\frac{(\ln\eta_{eq,n}+0.5\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i}^{m}))^{2}}{2\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i}^{m})}}\right).$ (54)
+$f_{H,n}(\eta_{eq,n}|\bm{d},\bm{\theta},\bm{\phi})$ $=\frac{\exp\left(\!{-\frac{(\ln\eta_{eq,n}+0.5\sigma_{eq,n}^{2}(\bm{d},\bm{\theta},\bm{\phi}))^{2}}{2\sigma_{eq,n}^{2}(\bm{d},\bm{\theta},\bm{\phi})}}\!\right)}{\eta_{eq,n}\sqrt{2\pi\sigma_{eq,n}^{2}(\bm{d},\bm{\theta},\bm{\phi})}}.$ (48)
+Then the unconditional PDF for the equivalent turbulent fading coefficient $\eta_{eq,n}$ can be obtained by averaging $\{\bm{d},\bm{\theta},\bm{\phi}\}$ on $\mathcal{F}$ out. Specifically, we have
+$f_{H,n}(\eta_{eq,n})$ $=\int_{\mathcal{F}}f(\bm{d},\bm{\theta},\bm{\phi}|\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F})$ (49)
+$\quad\quad\times f_{H,n}(\eta_{eq,n}|\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi},$
+where $f(\bm{d},\bm{\theta},\bm{\phi}|\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F})$ is given in (38). By submitting eqs. (38) and (39) into (49), we can obtain
+$f_{H,n}(\eta_{eq,n})$ (50)
+$\quad=\frac{\int_{\mathcal{F}}f(\bm{d},\bm{\theta},\bm{\phi})f_{H,n}(\eta_{eq,n}|\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\mathcal{F}}f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}$
+$\quad=\frac{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}f_{H,n}(\eta_{eq,n}|\bm{d},\bm{\theta},\bm{\phi})I_{n}f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}I_{n}f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}},$
+where we have used the equality $\int_{\mathcal{F}}g(\bm{d},\bm{\theta},\bm{\phi})f(\bm{d},\bm{\theta},\bm{\phi})$ $\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}=\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}I_{n}g(\bm{d},\bm{\theta},\bm{\phi})f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}$ for any function $g(\bm{d},\bm{\theta},\bm{\phi})$.
+This is because $I_{n}=0$ if $(\bm{d},\bm{\theta},\bm{\phi})$ is not in $\mathcal{F}$.
+Therefore, by substituting (48) into (50), we can finally obtain the unconditional PDF for the equivalent turbulent fading coefficient $\eta_{eq,n}$, which is given in (51) on the top of next page.
+$f_{H,n}(\eta_{eq,n})$ $=\frac{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}\frac{I_{n}}{\eta_{eq,n}\sqrt{2\pi\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i})}}\exp\left({-\frac{(\ln\eta_{eq,n}+0.5\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i}))^{2}}{2\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i})}}\right)f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}I_{n}f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}.$ (51)
+Then we can also adopt the MCI approach to estimate $f_{H,n}(\eta_{eq,n})$ by using the same sampling PDF given in (23).
+Using the same sampling points $\{\bm{s}_{n}^{1},\bm{s}_{n}^{2},\cdots,\bm{s}_{n}^{M}\}$, we can estimate the PDF of equivalent turbulent fading coefficient $f_{n}(\eta_{eq,n})$ as
+$f_{H,n}(\eta_{eq,n})$ $=\frac{1}{Count}\sum_{m=1}^{M}\frac{I_{n}(\bm{s}_{n}^{m})}{\eta_{eq,n}\sqrt{2\pi\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i}^{m})}}$ (52)
+$\times\exp\left({-\frac{(\ln\eta_{eq,n}+0.5\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i}^{m}))^{2}}{2\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i}^{m})}}\right).$
 According to (44), the PDF of the total equivalent turbulent fading coefficient $\eta_{eq}$ for $N$ scattering orders can be obtained as the convolution of the PDFs of $\{\frac{P_{1}}{P_{tot}}\eta_{eq,1},\frac{P_{2}}{P_{tot}}\eta_{eq,2},\cdots,\frac{P_{N}}{P_{tot}}\eta_{eq,N}\}$, i.e.,
-$f_{H}(\eta_{eq})$ $=\frac{P_{tot}}{P_{1}}f_{H,1}\left(\frac{P_{tot}}{P_{1}}\eta_{eq}\right)*\frac{P_{tot}}{P_{2}}f_{H,2}\left(\frac{P_{tot}}{P_{2}}\eta_{eq}\right)*\cdots*\frac{P_{tot}}{P_{N}}f_{H,N}\left(\frac{P_{tot}}{P_{N}}\eta_{eq}\right),$ (55)
+$f_{H}(\eta_{eq})$ $=\frac{P_{tot}}{P_{1}}f_{H,1}\left(\frac{P_{tot}}{P_{1}}\eta_{eq}\right)*\frac{P_{tot}}{P_{2}}f_{H,2}\left(\frac{P_{tot}}{P_{2}}\eta_{eq}\right)$ (53)
+$\quad*\cdots*\frac{P_{tot}}{P_{N}}f_{H,N}\left(\frac{P_{tot}}{P_{N}}\eta_{eq}\right),$
 where * denotes the convolution operation and $f_{H,n}(x)$ is calculated by (52).
-The pseudocode of the MCI approach is similar to the one we presented in [17].
+The pseudocode of the MCI approach is similar to the one we presented in [yuan2019monte].
 The only difference is that we need to add the estimation of turbulent fluctuation and distribution in each MCI process according to the equations given in (42) and (52), respectively.
-V Numerical Results and Discussions
-In this section, we present some numerical results to verify our theoretical analysis and discuss some interesting findings.
+We remark that though our derivation here is based on LN distribution for weak turbulence, we can also extend this derivation to moderate and strong turbulence cases.
+However, in moderate and strong turbulence, a large $C_{n}^{2}$ can result in a large Rytov variance $\sigma_{r}^{2}$ and thus a large turbulent variance $\text{Var}[\eta]\gg 1$ if LN distribution is adopted, which does not match practical moderate and strong turbulence with $\text{Var}[\eta]$ normally less than 2.
+An alternative way to extend our method to moderate and strong turbulence is using the Gamma-Gamma fading to calculate the turbulent variance $\text{Var}[\eta]$ since it never can never exceed 2; then we use this turbulent variance $\text{Var}[\eta]$ to calculate the parameter $\sigma_{I}^{2}$ in LN fading as $\sigma_{I}^{2}=\ln(\text{Var}[\eta]+1)$.
+Through this way we can estimate the distribution of turbulent fading coefficient for all turbulent conditions.
+V Simulation and Experimental Results
+In this section, we present some simulation and experimental results to verify our theoretical analysis.
 Unless otherwise specified, the simulation parameters we used in this section are listed in Table I.
@@
 $r$ $500$ m
-$[\theta_{T},\phi_{T}]$ $[45^{\circ},-90^{\circ}]$
-$[\theta_{R},\phi_{R}]$ $[45^{\circ},90^{\circ}]$
-$[\beta_{T},\beta_{R}]$ $[17^{\circ},30^{\circ}]$
+$\theta_{T},\phi_{T}] [$$45^{\circ},-90^{\circ}]$
+$\theta_{R},\phi_{R}] [$$45^{\circ},90^{\circ}]$
+$\beta_{T},\beta_{R}] [$$17^{\circ},30^{\circ}]$
 $A_{r}$ $1.77\times 10^{-4}{\text{m}}^{2}$
 $\lambda$ $2.6\times 10^{-7}$ m
-[…]
-$[k_{a}^{par},k_{s}^{ray},k_{s}^{mie}]$ $[0.802,0.266,0.284]$ ${\text{km}}^{-1}$
-$[\gamma,g,f]$ $[0.017,0.72,0.5]$
+$c$ $2.998\times 10^{8}$ $\text{m}/\text{s}$
+$k_{a}^{par},k_{s}^{ray},k_{s}^{mie}] [$0.802,0.266,0.284][0.802,0.266,0.284] ${\text{km}}^{-1}$
+$\gamma,g,f] [$0.017,0.72,0.5][0.017,0.72,0.5]
 $L_{0}$ $10^{2}$ m
@@
 $C_{n}^{2}$ $10^{-15}$ $\text{m}^{-2/3}$
-$\Delta t$ $0.02$ $\mu$s
+[…]
 $M$ $10^{6}$
-$N$ $3$
-V-A Results for Turbulence-Induced Scattering Effect
-Figure 2: Turbulence-induced scattering effect: (a) scattering coefficient $k_{s}^{tur}$ under different $C_{n}^{2}$ and $d_{0}$; (b) phase scattering function $p^{tur}(\theta_{s})$ at different $\theta_{s}$ and $d_{0}$
-We first explore the turbulence-induced scattering coefficient $k_{s}^{tur}$ and phase scattering function $p^{tur}(\theta_{s})$ under different turbulent parameters in Fig.
+NN 33
+V-A Simulation Results
+Figure 2: Phase scattering function due to turbulence in different $d_{0}$
+We first explore the turbulence-induced phase scattering function $p^{tur}(\theta_{s})$ under different scattering angle $\theta_{s}$ and correlation distance $d_{0}$ in Fig.
 2.
-Fig.
-2(a) presents the scattering coefficient $k_{s}^{tur}$ under different refractive-index structure parameter $C_{n}^{2}$ and correlation distance $d_{0}$.
-From Fig.
-2(a) we can see that the scattering coefficient $k_{s}^{tur}$ increases as either the turbulence strength or the correlation distance increases.
-Besides, for a given refractive-index structure parameter $C_{n}^{2}$, the slop of the turbulent scattering coefficient $k_{s}^{tur}$ for $d_{0}\gg\lambda$ is much smaller than that for $d_{0}\ll\lambda$, where $\lambda=2.6\times 10^{-7}$ m.
-These observations verified our theoretical analysis on Eq.
-(6).
-Fig.
-2(b) presents the phase scattering function $p^{tur}(\theta_{s})$ at different scattering angle $\theta_{s}$ and correlation distance $d_{0}$.
 Note that $L_{0}$ and $C_{n}^{2}$ have no impact on the phase scattering function.
 From Fig.
-2(b), we can see that, when $d_{0}\gg\lambda$, the scattering intensity rapidly increases as the scattering angle decreases.
-This indicates that the incident photons will have a high probability of maintaining their original traveling direction after scattering.
-By contrast, when $d_{0}\ll\lambda$, we can see that the phase scattering function varies slowly at different scattering angles and will finally approximate an isotropic scattering.
-These observations verified our theoretical analysis on Eq.
-(7).
-The impacts of turbulence-induced scattering effect on the receiving power of NLOS UV communications should be explored by considering both the scattering coefficient $k_{s}^{tur}$ and the phase scattering function $p^{tur}(\theta_{s})$.
-Combining the results in Figs.
-2(a) and 2(b), we have the following observations:
-  • When $d_{0}\ll\lambda$, we have $k_{s}^{tur}\to 0$; in this case, though the phase scattering function approaches an isotropic scattering, the overall impact of the turbulence-induced scattering effect is negligible compared with the scattering effect due to particles; therefore, we can always ignore the turbulence-induced scattering effect when $d_{0}\ll\lambda$.
-  • When $d_{0}\gg\lambda$, the scattering is almost a complete forward scattering since almost all the scattering intensity concentrates at $\theta_{s}=0^{\circ}$; in this case, though there exists large scattering coefficient $k_{s}^{tur}\gg 1$, the overall turbulence-induced scattering can be equivalently regarded as a rectilinear propagation with negligible divergence angle, which means no evident impact on the final receiving power can be observed; therefore, we can also ignore the turbulence-induced scattering effect when $d_{0}\gg\lambda$.
-  • When $d_{0}\sim\lambda$, the phase scattering function is not a complete forward scattering; then we cannot ignore the turbulence-induced scattering effect if the scattering coefficient $k_{s}^{tur}$ is not negligible compared with $k_{s}^{par}$, which can happen under strong turbulent conditions; for example, when $d_{0}=10^{-7}$ m, $C_{n}^{2}=10^{-13}$ $\text{m}^{-2/3}$, and $L_{0}=10^{3}$ m, we have $k_{s}^{tur}=0.586$ $\text{km}^{-1}$, which is comparable with $k_{s}^{par}=0.550$ $\text{km}^{-1}$.
-Figure 3: CIR results under different outer scale parameters $L_{0}$: (a) $r=200$ m; (b) $r=500$ m; (c) $r=800$ m
-Let us explore the cases with $d_{0}\sim\lambda$.
-Fig.
-3 presents the CIR results under different outer scale parameters $L_{0}$ in different communication distances $r\in\{200\ \text{m},500\ \text{m},800\ \text{m}\}$ with $d_{0}=10^{-7}$ m and $C_{n}^{2}=10^{-13}$ $\text{m}^{-2/3}$.
-Comparing Figs.
-3(a), 3(b), and 3(c), we can see that when the communication distance is small, the turbulence-induced scattering can enhance the receiving power.
-However, as the communication distance increases, the enhancement due to the turbulence-induced scattering gradually decreases and finally the turbulence-induced scattering will introduce an extra attenuation on receiving power in large communication distance.
-Here we can explain this distance-dependent scaling effect on the receiving power.
-Since the single-scattering is the major scattering here, we roughly consider the single-scattering only.
-As we know, an increase in the scattering coefficient can not only increase the scattering strength but also increase the extinction strength and shorten the photon traveling distance.
-The expectation of the photon traveling distance can be estimated as $1/k_{e}^{tot}$ and here we have $1/k_{e}^{tot}=1/(k_{s}^{par}+k_{a}^{par})\approx 739.6$ m by ignoring the turbulence-induced scattering.
-Then when the communication distance is smaller than $739.6$ m, an increase in the scattering coefficient will shorten the expected traveling distance and thus increase the scattering strength in the common volume, and vice versa.
-Therefore, when $r=200$ m, the turbulence-induced scattering mainly acts as a scattering effect enhancer; whereas when $r=800$ m, the turbulence-induced scattering mainly acts as an extinction effect enhancer.
-Though we have demonstrated that the turbulence-induced scattering effect can impact the final receiving power when $d_{0}\sim\lambda$, in practical implementations, we have $d_{0}\gg\lambda$ almost always holds.
-For example, the correlation distance $d_{0}$ is normally in between $l_{0}$ and $L_{0}$, where $l_{0}$ is the inner scale parameter representing the smallest turbulent eddy size; and $l_{0}$ is at the same order of magnitude with the Fresnel zone $R_{F}=(L/k)^{1/2}$, where $k$ is the light wave number and $L$ is the light propagating distance.
-For typical UV communication scenarios, we have $10^{0}\ \text{m}\leq L\leq 10^{4}\ \text{m}$ and $2.034\times 10^{-4}\ \text{m}\leq R_{F}\leq 2.034\times 10^{-2}\ \text{m}$, which results in $d_{0}\sim R_{F}\gg\lambda$.
-Therefore, we can conclude that the turbulence-induced scattering effect can always be ignored in typical UV communication scenarios.
-V-B Results for Turbulent Fluctuation Effect
-Figure 4: Turbulent fluctuations under different number of sampling points
+2, we can see that, the scattering intensity rapidly increases as the scattering angle decreases.
+This indicates that the incident photons almost always maintain their original traveling direction after scattering since almost all the scattering intensity concentrates at $\theta_{s}=0^{\circ}$.
+In this case, the overall turbulence-induced scattering can be equivalently regarded as a rectilinear propagation with negligible divergence angle, which means no evident impact on the final receiving power can be observed.
+Therefore, we can ignore the turbulence-induced scattering effect in UV wavebands.
+Figure 3: Turbulent fluctuations under different number of sampling points: (a) conventional MC method in [wang2013characteristics]; (b) improved MCI method
 Then we explore the turbulent fluctuation effects.
@@
 3 presents the estimated turbulent fluctuation under different number of sampling points $M$.
-As we can see from Fig.
-4, the estimated turbulent fluctuation $\sigma^{2}_{tur}$ or $\sigma_{tur,n}^{2}$ will become stable as $M$ increases.
-Besides, we can also roughly estimate the turbulent variance by using the method proposed for the single-scattering link in [47].
+To numerically demonstrate the necessity of our model, we first use the conventional Monte-Carlo method used in [wang2013characteristics].
+The simulation results, obtained under varying $M$ with 100 realizations, are presented in Fig.
+3(a).
+Consistent with the prediction from Eq.
+(31), the turbulent fluctuation estimated by the conventional Monte-Carlo method exhibits a decrease as $M$ increases from $10^{2}$ to $10^{5}$.
+The dependency on $M$ indicates that the results are dominated by Monte-Carlo sampling variance rather than physical turbulence.
+In contrast, our improved MCI method provides stable values for turbulent variance, as shown in Fig.
+3(b), where the estimated turbulent fluctuation $\sigma^{2}_{tur}$ or $\sigma_{tur,n}^{2}$ becomes stable as $M$ increases.
+Besides, we can also roughly estimate the turbulent variance by using the method proposed for the single-scattering link in [ding2011turbulence].
 Specifically, the turbulent variance can be approximated as $\sigma_{tur,1}^{2}\approx\exp(\sigma_{I}^{2}(r_{1})+\sigma_{I}^{2}(r_{2}))-1\approx 0.0493$, which is quite close to our result $\sigma_{tur,1}^{2}=0.0492$ in Fig.
+3.
+This verified that our proposed estimation method for turbulent fluctuation can well estimate the turbulent variance.
+Figure 4: Turbulent fluctuations under various system geometries: (a) turbulent fluctuations under different communication distances; (b) turbulent fluctuations under different zenith angles
+Then we present the turbulent fluctuations results under various system geometries in Fig.
 4.
-This verified that our proposed estimation method for turbulent fluctuation can well estimate the turbulent variance.
-Figure 5: Turbulent fluctuations under various system geometries: (a) turbulent fluctuations under different communication distances; (b) turbulent fluctuations under different zenith angles; (c) turbulent fluctuations under different divergence angles
-Then we present the turbulent fluctuations results under various system geometries in Fig.
+From Figs.
+4(a) and 4(b), we can see that the turbulent variance increases as either the communication distance $r$ increases or the zenith angle $\theta_{T}$ decreases  [^1] This indicates that a worse bit-error rate (BER) performance will be obtained when the communication distance or the elevation angle increases, which is compatible with the results in [he2024demonstration]., which is compatible with reported experimental results in [chen2014experimental, liao2014turbulence].
+Besides, we can also observe that the turbulence variance will approach a stable value for strong turbulent conditions in Figs.
+4(a) and 4(b).
+This is because the channel will become a saturated turbulent channel under strong turbulence as either the communication distance $r$ increases or the zenith angle $\theta_{T}$ decreases, which is compatible with the saturated turbulence case with $\sigma_{r}^{2}\to\infty$.
+Figure 5: Comparison of the proposed model with existing related works [ding2011turbulence] and [chen2014experimental]: (a) Comparison of single-scattering turbulent fluctuation with [ding2011turbulence]; (b) Comparison of total turbulent fluctuation with experimental results in [chen2014experimental].
+To further validate the accuracy of the proposed model, we compare our results with the single-scattering turbulence model in [ding2011turbulence] and the experimental data reported in [chen2014experimental], respectively, as shown in Fig.
 5.
+The simulation parameters are set according to the experimental configuration in [chen2014experimental], with the refractive-index structure parameter remains $C_{n}^{2}=10^{-15}\text{m}^{-2/3}$ and $\theta_{R}=40^{\circ}$.
+Two specific geometric configurations were selected for this comparison: (i) $r=1000$ m with $\theta_{T}=40^{\circ}$, and (ii) $r=1500$ m with $\theta_{T}=60^{\circ}$.
+Fig.
+5(a) compares the single-scattering turbulent fluctuation $\sigma^{2}_{tur,1}$ calculated by our model against the analytical model in [ding2011turbulence].
+It can be observed that our results highly consist with the existing analytical model across different geometric configurations.
+Furthermore, Fig.
+5(b) compares the total turbulent fluctuation $\sigma^{2}_{tur}$ of our multiple-scattering model with the experimental measurements from [chen2014experimental].
+The results indicate that our model fits the experimental data well.
+Figure 6: Distribution of turbulent fading coefficient under different turbulent conditions with $r=200$ m and $r=500$ m: (a) weak turbulence; (b) moderate turbulence; (c) strong turbulence; (d) PDF comparison with [liao2014turbulence]
+Then we present the estimated PDF for the turbulent fading coefficients under different communication distances $r=200$ m and $r=500$ m with different turbulent conditions in Fig.
+6.
 From Figs.
-5(a), 5(b) and 5(c), we can see that the turbulent variance increases as either the communication distance $r$, the elevation angle $\pi/2-\theta_{T}$, or the divergence angle $\beta_{T}$ increases, which is compatible with reported experimental results in [53, 54].
-Besides, we can also observe that the turbulence variance will approach a stable value for strong turbulent conditions in Figs.
-5(a), 5(b) and 5(c).
-This is because the channel will become a saturated turbulent channel under strong turbulence as either the communication distance $r$, the elevation angle $\pi/2-\theta_{T}$, or the divergence angle $\beta_{T}$ increases, which is compatible with the saturated turbulence case with $\sigma_{r}^{2}\to\infty$ at the end of Section II.
-Figure 6: Distribution of turbulent fading coefficient under different distances: (a) $r=200$ m; (b) $r=500$ m; (c) $r=800$ m
-At last, we present the estimated probability density for the turbulent fading coefficients under different communication distances $r\in\{200\ \text{m},500\ \text{m},800\ \text{m}\}$ with $C_{n}^{2}=10^{-17}$ $\text{m}^{-2/3}$ in Fig.
-6.
-The simulation results of $f_{H,n}(\eta_{eq,n})$ for the $n$th order scattering are fitted with a weighted summation of $n$ Gaussian distributions, and the simulation results of $f_{H}(\eta_{eq})$ are fitted with a Gaussian distribution.
+6(a) to 6(c), we can observe that the simulation results of $f_{H}(\eta_{eq})$ can be well approximated by a LN distribution in all turbulent conditions.
+Specifically, when the turbulence is weak, a Gaussian distribution can also well fit the obtained PDF, which is reasonable because an LN distribution will approach a Gaussian distribution when the variance is small.
+Besides, we also compare our model with the turbulent distribution reported in an experimental work [liao2014turbulence] in Fig.
+6(d).
 From Fig.
-6, we can see that the Gaussian fitting can well match the simulation results.
-Here we give an explanation on the single-scattering only.
-For a single-scattering, we can divide the common volume $V$ into many small volumes $\{v_{1},v_{2},\cdots,v_{L}\}$.
-Then according to the LN assumption in [47], the receiving power $P_{r}^{l}$ contributed from the $l$th volume $v_{l}$ follows a LN distribution.
-Then according to the central limit theorem, the total receiving power $P_{r}=\sum_{l=1}^{L}P_{r}^{l}$ will converge to a Gaussian distribution when $L$ is large.
+6(d) we can see that our model can well approximate the experimental results reported in [liao2014turbulence].
+V-B Experimental Results
+Figure 7: Experimental platform and typical receiving voltage under different zenith angles: (a) experimental platform; (b) $r=50$ m and $\theta_{T}=\theta_{R}=50^{\circ}$; (c) $r=50$ m and $\theta_{T}=\theta_{R}=60^{\circ}$
+We established a experimental platform for measuring the turbulent fluctuation, shown in Fig.
+7(a).
+At the transmitter end, a signal generator is employed to generater on-off keying signals and the signal is directed to a LED driver.
+Then the LED driver drives the UV LED array and transmitting UV signals.
+The LED array consists of 8 LEDs with transmitting light power up to 30 mW and a divergence angle around $10^{\circ}$ (half-power divergence angle of Lambertian light source) for each LED.
+At the receiver end, a photomultiplier tube (PMT) combined with a UV filter is used to detect the scattered UV signal, then the detected signal is amplified by a trans-impedance amplifier and directed to an oscilloscope.
+The receiving FOV angle of the PMT is around $45^{\circ}$.
+The oscilloscope records the received signal and saves it for further processing.
+Two goniometers are employed to measure the transmitting and receiving zenith angles and two portable power suppliers are employed for power supply.
+The experiments were conducted on April 29, 2025 from 5:00 pm to 7:00 pm at Haidian campus of Beijing university of posts and telecommunications with ground wind speed around 4.5 m/s.
+The measurements were carried at two distances $r\in\{50\text{m},80\text{m}\}$ and six zenith angles with $\theta_{T}=\theta_{R}\in\{20^{\circ},30^{\circ},40^{\circ},50^{\circ},60^{\circ},70^{\circ}\}$.
+Two typical receiving signals in voltage under $(r=50\text{m},\theta_{T}=\theta_{R}=50^{\circ})$ and $(r=50\text{m},\theta_{T}=\theta_{R}=60^{\circ})$ are shown in Figs.
+7(b) and 7(c), respectively.
+Figure 8: Experimental results of average receiving voltages and scintillation indices under different zenith angles: (a) Average receiving voltage; (b) Scintillation index
+Then we calculate the average receiving voltage and corresponding scintillation index under different transmitting distances and zenith angles.
+The scintillation index is calculated as $\sigma_{I}^{2}=\frac{\text{E}[R_{V}^{2}]}{\text{E}^{2}[R_{V}]}-1$, where $R_{V}$ is the receiving voltage.
+The obtained average voltage and scintillation index are shown in Figs.
+8(a) and 8(b), respectively.
+From Fig.
+8(a) we can observe that a larger zenith angle corresponds to a larger average receiving voltage.
+Besides, a larger communication distance generally corresponds a smaller average receiving voltage.
+From Fig.
+8(b), we can see that the scintillation index decreases as the zenith angle increases.
+Besides, the scintillation index generally increases as the communication distance increases.
+These experimental results verified the simulation results in Fig.
+4.
+We also notice that the obtained data points demonstrated mess results when the zenith angle is too large.
+This is because when the zenith angle is too large, there exits small line-of-sight signals, which may affect the receiving signals.
+Figure 9: Experimental data and corresponding Gaussian fitting results under different distances: (a) $r=50$ m; (b) $r=80$ m
+At last, we present the obtained distribution results from the normalized experimental data and the corresponding Gaussian fitting results under different communication distances and zenith angles in Fig.
+9.
+From Fig.
+9 we can see that though the data points is not enough, we can also observe that a Gaussian fitting can fit the distribution of receiving signals under different communication distances and zenith angles.
+Because our experiments are conducted in relatively short range, a weak turbulence condition can be assumed, which verified our simulation results in Fig.
+6(a).
 VI Conclusion
-Existing works on the modeling of multiple-scattering turbulent channels either ignored the turbulence-induced scattering effect or erroneously estimated the turbulent fluctuation effect on the NLOS UV communications.
-In this paper, we established a multiple-scattering turbulent channel model considering both the turbulence-induced scattering effect and the turbulent fluctuation effect.
-Numerical results demonstrated that, for typical UV communication scenarios, the turbulence-induced scattering effect can always be ignored; and the turbulent fluctuation will increase as either the communication distance, the elevation angle, or the divergence angle increases, which is compatible with existing experimental results.
-Besides, we also demonstrated that the distribution of the overall turbulent fading coefficient for the multiple-scattering process can be approximated as a Gaussian distribution.
+Existing works on the modeling of turbulent channels for NLOS UV communications either focused on the single-scattering cases or estimated the turbulent fluctuation effect in an unreliable way based on an MCS approach.
+In this paper, we established a turbulent multiple-scattering channel model by using an MCI approach, which is more efficient and interpretable compared with the MCS approach.
+Numerical results demonstrated that, for typical UV communication scenarios, the turbulence-induced scattering effect can always be ignored; and the turbulent fluctuation will increase as either the communication distance increases or the zenith angle decreases, which is compatible with existing experimental results and also with our experimental results.
+Besides, we demonstrated numerically that the distribution of the overall turbulent fading coefficient for the multiple-scattering process can be approximated as LN distributions under various turbulent conditions; and we also demonstrated both numerically and experimentally that the turbulent fading can be approximated as a Gaussian distribution under weak turbulence.
+Our work can help researchers to accurately estimate communication performance, e.g., BER and outage probability, in turbulent channels, which is important for designing their UV communication system components, e.g., modulation schemes and link budget.
+For example, according to our findings, a more robust modulation scheme such as pulse-position modulation scheme rather than pulse-amplitude modulation scheme is preferred in large communication distances or elevation angles to combat the large turbulent fading.
+Besides, a higher risk of link interruption happens in a larger communication distance or elevation angle, and therefore requires a larger link budget in designing the communication system.
+We have to point out that our experimental validation was conducted in an urban environment under relatively weak turbulence conditions.
+Turbulence characteristics can vary significantly across different scenarios due to variations in temperature gradients, humidity levels, and atmospheric composition.
+Though the effects of various weather conditions on UV NLOS communication was investigated in one of our previous work [shan2023non], the combination between the turbulent effects and the weather conditions has not been studied yet.
+In future work, we will further explore turbulence effects in diverse potential UV NLOS communication applications, such as maritime conditions with high humidity, desert regions with strong thermal turbulence, and high-altitude scenarios with reduced atmospheric density.
+Investigating how these environmental variations influence turbulence-induced fading could provide a more comprehensive understanding of UV NLOS channel characteristics across a broader range of channel conditions.